### v2

# Gradient Methods with Online Scaling Part II. Practical Aspects

Abstract
Part I of this work [14] establishes online scaled gradient methods (OSGM) [^1] This paper extends two previous works [13] and [5]., a framework that utilizes online convex optimization to adapt stepsizes in gradient methods. This paper focuses on the practical aspects of OSGM. We leverage the OSGM framework to design new adaptive first-order methods and provide insights into their empirical behavior. The resulting method, OSGM-Best, matches the performance of quasi-Newton variants while requiring less memory and cheaper iterations. We also extend OSGM to nonconvex optimization and outline directions that connect OSGM to existing branches of optimization theory and practice.
1 Introduction
Consider gradient descent applied to a smooth convex problem $f^{\star}\coloneqq\min_{x\in\mathbb{R}^{n}}f(x)$:
$x^{k+1}=x^{k}-P_{k}\nabla f(x^{k}),$ (1)
where $P_{k}\in\mathbb{R}^{n\times n}$ is a matrix stepsize. The first part of this work [14] introduces OSGM, a framework that uses online learning to adjust stepsize $\{P_{k}\}$ on the fly and achieves problem-dependent acceleration.
We briefly summarize the idea of OSGM introduced in Part I [14]. The analyses of algorithms for smooth optimization often show that each iteration yields some improvement. For instance, in linearly convergent methods, a common improvement is a contraction of the form $\tfrac{f(x^{k+1})-f^{\star}}{f(x^{k})-f^{\star}}\leq 1-\tfrac{1}{\kappa}$. Here $\kappa\geq 1$ often denotes some condition number. If this contraction ratio bound holds uniformly for all $k$, chaining the per iteration improvements over $K$ steps yields the rate
$\tfrac{f(x^{K+1})-f^{\star}}{f(x^{1})-f^{\star}}=\textstyle\prod_{k=1}^{K}\tfrac{f(x^{k+1})-f^{\star}}{f(x^{k})-f^{\star}}\leq(1-\tfrac{1}{\kappa})^{K}.$
Many proofs first establish improvement at every iteration and then chain these improvements to prove convergence. However, this approach requires monotonic improvement. OSGM instead chains the contraction first:
$\tfrac{f(x^{K+1})-f^{\star}}{f(x^{1})-f^{\star}}=\textstyle\prod_{k=1}^{K}\tfrac{f(x^{k+1})-f^{\star}}{f(x^{k})-f^{\star}}\leq(\tfrac{1}{K}\textstyle\sum_{k=1}^{K}\tfrac{f(x^{k+1})-f^{\star}}{f(x^{k})-f^{\star}})^{K}.$ (2)
To guarantee convergence, it suffices to have the average contraction ratio $\tfrac{1}{K}\textstyle\sum_{k=1}^{K}\tfrac{f(x^{k+1})-f^{\star}}{f(x^{k})-f^{\star}}$ be sufficiently small. In other words, convergence is possible even if some iterations worsen the objective. When the iterates $\{x^{k}\}$ are generated via the update (1), the contraction ratio attributed to one iteration, starting from $x$, can be written as $r_{x}(P)=\tfrac{f(x-P\nabla f(x))-f^{\star}}{f(x)-f^{\star}}$, which depends on the stepsize matrix. In view of (2), it suffices to select $\{P_{k}\}$ sequentially to minimize $\tfrac{1}{K}\textstyle\sum_{k=1}^{K}r_{x^{k}}(P_{k})$, the average contraction ratio. This perspective formulates stepsize selection as a sequential decision-making problem, a central topic in online learning [27, 18]. If $\{P_{k}\}$ are chosen by a no-regret online learning algorithm, we obtain the guarantee: $\tfrac{1}{K}\textstyle\sum_{k=1}^{K}r_{x^{k}}(P_{k})\leq\tfrac{1}{K}\textstyle\sum_{k=1}^{K}r_{x^{k}}(\hat{P})+o(1)$ for any benchmark stepsize $\hat{P}$, including the optimal preconditioner [15, 31]. The relation further yields the convergence guarantee
$\tfrac{f(x^{K+1})-f^{\star}}{f(x^{1})-f^{\star}}\leq(\tfrac{1}{K}\textstyle\sum_{k=1}^{K}r_{x^{k}}(\hat{P})+o(1))^{K}\text{~~~for any $\hat{P}$.}$
More generally, OSGM models stepsize selection in a gradient-based method as a sequential decision-making problem between a stepsize scheduling agent and a landscape agent. At each iteration:
  1. 1.
    Scheduler makes decision $P_{k}$ and proposes an update $x^{k+1/2}=x^{k}-P_{k}\nabla f(x^{k})$.
  2. 2.
    Landscape chooses the next iterate $x^{k+1}=\mathcal{M}(x^{k},x^{k+1/2})$ and provides feedback $\ell_{x^{k}}(P_{k})$ to the scheduler.
  3. 3.
    Scheduler updates the stepsize using an online learning algorithm $\mathcal{A}(P_{k},\{\ell_{x^{j}}\}_{j\leq k})$.
Different instantiations of $(\ell_{x},\mathcal{M},\mathcal{A})$ yield different algorithm variants, each with its own convergence properties. Part I of this work develops the theoretical foundations of OSGM, providing specific instantiations that achieve both global and local convergence guarantees.
1.1 Contributions
Building on the theoretical foundations established in Part I, the second part focuses on the practical aspects of OSGM and outlines several future directions.
Insights into algorithm behavior.
While Part I focuses on complexity analysis, this work investigates the empirical behavior of OSGM, with a particular focus on OSGM-H [14]. We develop a more detailed analysis that explains the empirical behavior of OSGM-H and other hypergradient-type methods [2]. These insights lead to concrete algorithmic recommendations.
A practical variant of OSGM.
Part I developed OSGM variants based on gradient descent. This paper shows how to use the same ideas to choose stepsizes in other gradient-based methods, such as heavy-ball momentum. The strategies above, when combined with heavy-ball momentum, give rise to OSGM-Best: a robust variant of OSGM that performs competitively with L-BFGS [25] while requiring less memory and cheaper iterations.
Nonconvex optimization.
Although OSGM is grounded in online convex optimization, we demonstrate that the analysis can be extended to smooth nonconvex problems by introducing regularization in the stepsize space. Under the Polyak-Łojasiewicz (PL) condition, the resulting OSGM variants provably adapt to locally convex regions of the landscape. Experiments on standard nonconvex benchmarks support these theoretical claims.
Connection to other optimization techniques.
We highlight connections between OSGM and a range of ideas in optimization, including the Barzilai-Borwein step [1], accelerated gradient descent [8], proximal gradient methods [3], and the performance estimation problem (PEP) [35]. These connections highlight the potential of OSGM as a valuable addition to the modern optimization toolbox.
1.2 Notations
We use $\|\cdot\|$ to denote the Euclidean norm. Letters $A,a$ denote matrices and scalars. $\|A\|_{F}\coloneqq\sqrt{\sum_{ij}a_{ij}^{2}}$ denotes the matrix Frobenius norm. We use $\langle\cdot,\cdot\rangle$ for Euclidean or Frobenius inner product. $\Pi_{\mathcal{P}}[\cdot]$ denotes the orthogonal projection onto a closed convex set $\mathcal{P}$. We use $\mathcal{X}^{\star}=\{x:f(x)=f^{\star}\}$ to denote the optimal set of $f$ and $\mathcal{S}^{\star}$ to denote the set of stationary points of $f$; $\operatorname{dist}(P,\mathcal{P})\coloneqq\|P-\Pi_{\mathcal{P}}[P]\|_{F}$ denotes the distance between a point $P$ and a closed convex set $\mathcal{P}$; $\operatorname{diam}(\mathcal{P})=\max_{X,Y\in\mathcal{P}}\|X-Y\|_{F}$ denotes the diameter of set $\mathcal{P}$ in Frobenius norm. We use superscript $x^{k}$ to index algorithm iterates and subscript $P_{k}$ to index the stepsize sequence. A function $f$ satisfies $\mu$-PL condition if $f(x)-f^{\star}\leq\frac{1}{2\mu}\|\nabla f(x)\|^{2}$. To emphasize that the feedback functions have stepsize $P$ as their argument, we will use $h_{x}(P)$ instead of $h_{x}$ in most contexts.
Structure of the paper.
This paper is organized as follows. Section˜2 provides insights into the convergence behavior of OSGM-H observed in the literature. These insights, combined with a heavy-ball momentum enhancement, constitute OSGM-Best, a variant that yields competitive practical performance. Section˜3 generalizes OSGM to smooth nonconvex optimization. Section˜4 presents numerical experiments on both convex and nonconvex optimization problems. Section˜5 outlines several future directions and shows the connection between OSGM and existing optimization theory.
2 Algorithmic behavior and ingredients of practical OSGM
In this section, we motivate the key ingredients of OSGM-Best that improve the practical convergence of OSGM. We provide insights into each ingredient through either rigorous analysis or illustrative examples. Our analysis focuses on OSGM-H (Algorithm˜1), OSGM using hypergradient feedback, although most of the insights extend naturally to other variants of OSGM. The hypergradient feedback [14] takes the form
$h_{x}(P)=\tfrac{f(x-P\nabla f(x))-f(x)}{\|\nabla f(x)\|^{2}}.$
1input: $x^{1}$, $P_{1}\in\mathbb{R}^{n\times n}$, online gradient stepsize $\eta>0$
2 for $k=1,2,\dots$ do
3 $x^{k+1/2}=x^{k}-P_{k}\nabla f(x^{k})$
4   Let $x^{k+1}$ satisfy $f(x^{k+1})\leq\min\{f(x^{k+1/2}),f(x^{k})\}$
5 $P_{k+1}=P_{k}-\eta\nabla h_{x^{k}}(P_{k})$
6 end for
Algorithm 1 Monotone OSGM-H
1input: $x^{1}$, $P_{1}\in\mathcal{P}=\mathbb{R}^{n\times n}$, $\eta>0$
2 for $k=1,2,\dots$ do
3 $P_{k+1}=P_{k}-\eta\nabla h_{x^{k}}(P_{k})$
4 $x^{k+1}=x^{k}-P_{k+1}\nabla f(x^{k})$
5 end for
Algorithm 2 Classic-HDM
The feedback $h_{x}(P)$ is motivated by the classical descent lemma $f(x-\frac{1}{L}\nabla f(x))\leq-\frac{1}{2L}\|\nabla f(x)\|^{2}$ and is closely related to the classic hypergradient descent heuristic (Algorithm˜2) studied in the literature [2, 32]. However, there are two key differences: the monotonicity condition imposed on Algorithm˜1 and the order in which $P_{k}$ and $x^{k}$ are updated. These seemingly subtle distinctions turn out to be critical for an effective variant of OSGM-H in both theory and in practice. In this section, we detail our design and address the following aspects:
  • Monotone landscape action for OSGM-H. Is it necessary to enforce monotonicity in OSGM-H?
  • Limitations of hypergradient descent in the literature. How does the order of the updates in OSGM-H matter?
  • Heavy-ball momentum and modified feedback. Can momentum accelerate the convergence of OSGM-H?
After discussing each ingredient in turn, we combine them into one recipe: OSGM-Best.
2.1 Monotone landscape action for hypergradient feedback
The convergence analysis for OSGM-H in Part I [14] always requires a monotone landscape action to ensure at least $f(x^{k+1})\leq f(x^{k})$. In practice, OSGM-H is still observed to converge without enforcing monotonicity, but it exhibits periodic “spikes” in the objective (Figure˜1 (left)). Similar spiky patterns have also been observed in other methods, such as heavy-ball momentum [7] and the Barzilai-Borwein method [11].
Figure 1: Spiky behavior of Vanilla OSGM-H on a quadratic function.
In this section, we show that this requirement is not an artifact of analysis. Actually, one can construct an example for which the spike of OSGM-H can be arbitrarily high and can happen arbitrarily late. This precludes the possibility of establishing a standard convergence rate in terms of $K$ without the monotone landscape action.
Example.
Consider a simple 2D quadratic function parametrized by $\kappa\gg 1$:
$\min_{x=(x_{1},x_{2})\in\mathbb{R}^{2}}f(x)=\tfrac{1}{2}x_{1}^{2}+\tfrac{\kappa}{2}x^{2}_{2}.$
The objective $f$ is $\kappa$-smooth and $1$-strongly convex with condition number $\kappa$ and $x^{\star}=(0,0)$. The minimax optimal stepsize for this problem is $\alpha^{\star}=\frac{2}{1+\kappa}\approx\frac{2}{\kappa}$, and stepsizes larger than $\alpha^{\star}$ can lead to divergence. Scalar version of OSGM-H (without enforcing monotonicity) reads
$\left(\begin{smallmatrix}x^{k+1}_{1}\\ x^{k+1}_{2}\end{smallmatrix}\right)$ $=\left(\begin{smallmatrix}(1-\alpha_{k})x^{k}_{1}\\ (1-\kappa\alpha_{k})x^{k}_{2}\end{smallmatrix}\right),$
$\alpha_{k+1}$ $=\alpha_{k}+\eta\Big[\underbrace{\tfrac{(x^{k}_{1})^{2}}{(x^{k}_{1})^{2}+\kappa^{2}(x^{k}_{2})^{2}}(1-\alpha_{k})+\tfrac{\kappa^{2}(x^{k}_{2})^{2}}{(x^{k}_{1})^{2}+\kappa^{2}(x^{k}_{2})^{2}}(1-\kappa\alpha_{k})}_{\Delta_{k}}\Big].$ (3)
The sign of stepsize change $\Delta_{k}$ is determined by both $1-\alpha_{k}$ and $1-\kappa\alpha_{k}$. When $\alpha_{k}>\frac{1}{\kappa}$, the value $1-\kappa\alpha_{k}<0$ from the second coordinate will try reducing the stepsize, while $1-\alpha_{k}$ from the first coordinate tends to increase the stepsize until $\alpha_{k}\geq 1\gg\alpha^{\star}$. However, these two terms are weighted by quantities that depend on the location of $x^{k}$ on the landscape. In particular, if $x^{k}_{1}\gg\kappa x^{k}_{2}$, then $\tfrac{\kappa^{2}(x^{k}_{2})^{2}}{(x^{k}_{1})^{2}+\kappa^{2}(x^{k}_{2})^{2}}$ is close to 0, and the value $1-\kappa\alpha_{k}$ from the second coordinate will be disregarded. It can therefore lead to overly large stepsize and divergence. Suppose OSGM-H is initialized such that $x^{k}_{1}\gg\kappa x^{k}_{2}$. Then each spike in Figure˜1 can be decomposed into three phases.
Myopic stepsize increase. When $x^{k}_{1}\gg\kappa x^{k}_{2}$, $\Delta_{k}\approx 1-\alpha_{k}$, and OSGM-H drives the stepsize to $1$ to accelerate the convergence in the direction of $x_{1}$. However, the magnitude of $x_{2}$ explodes since $|1-\kappa\alpha_{k}|\gg 1$.
Divergence and recovery. Though the divergence of $x_{2}$ starts from a small initial value, it finally becomes large enough that the second term $\tfrac{\kappa^{2}(x^{k}_{2})^{2}}{(x^{k}_{1})^{2}+\kappa^{2}(x^{k}_{2})^{2}}$ in (3) dominates, decreasing $\alpha_{k}$. However, OSGM-H requires many iterations to decrease $\alpha_{k}$ to $\tfrac{2}{\kappa}$, and $x^{k}_{2}$ continues to increase during this phase, resulting in a spike in the objective.
Periodic spiky behavior. After OSGM-H decreases the stepsize below $\tfrac{2}{\kappa}$, the second coordinate $x^{k}_{2}$ decreases sharply. However, $x^{k}_{2}$ decreases much faster than $x^{k}_{1}$ does since $1-\kappa\alpha_{k}\ll 1-\alpha_{k}$ if $\alpha_{k}=\mathcal{O}(\tfrac{1}{\kappa})$. Then OSGM-H returns to the state $x_{1}^{k}\gg\kappa x_{2}^{k}$, and repeats the same divergence-recovery loop. Notice that the magnitude $|x_{1}|$ consistently decreases (Figure˜1 (right)), since $\alpha_{k}$ remains less than $1$ and thus $|1-\alpha_{k}|<1$. Hence, the algorithm ultimately converges after several loops. But this analysis also precludes a standard convergence proof, as we can pick $\delta\rightarrow 0$ to construct a bad starting point $(1,\delta)$ for any (sufficiently large) iteration $K$ so that a spike happens around $K$ and $f(x^{K})>f(x^{1})$.
To quantify the spiky behavior, suppose OSGM-H starts from $(x^{1}_{1},x^{1}_{2})=(1,\delta)$ for some $\delta>0$ to be determined later, and the stepsize is initialized by $\alpha_{1}=0$. Lemma˜2.1 shows that OSGM-H increases stepsize by at least a constant factor of $\eta\leq\tfrac{1}{\kappa}$ when $\alpha_{k}$ is small enough and $x^{k}_{1}\gg\kappa x^{k}_{2}$.
Lemma 2.1.
Let $\kappa\geq 2$. If $\alpha_{k}\leq\tfrac{1}{2}$ and $|x^{k}_{1}|\geq\sqrt{2}\kappa^{3/2}|x^{k}_{2}|$, then $\alpha_{k+1}\geq\alpha_{k}+\tfrac{\eta}{4}$.
Assume $\kappa$ is large enough to have a constant $a$ such that $\tfrac{2}{\kappa}\ll\tfrac{a}{2}<\tfrac{1}{2}$, say $a=\tfrac{1}{3}$. Note that if the condition in Lemma˜2.1 is always satisfied, OSGM-H takes at most $\lceil\tfrac{4a}{\eta}\rceil$ iterations to increase the stepsize from $\alpha_{1}=0$ to at least $a$. During the stepsize increasing period, the magnitude $|x^{k}_{2}|$ is at most expanded by a factor $|1-\kappa\alpha_{k}|<\kappa-1$ since $\alpha_{k}<1$. Therefore, we can choose $\delta>0$ small enough such that the iterates $\{(x^{k}_{1},x^{k}_{2})\}_{k\leq K}$ always satisfy the condition in Lemma˜2.1 but $(x^{K+1}_{1},x^{K+1}_{2})$ violates the same condition. Then OSGM-H starts to decrease the stepsize after iteration $K+1$, but the stepsize $\alpha_{K+1}>a\gg\tfrac{2}{\kappa}$ and thus $x_{2}$ continues to expand, causing a spike in the objective. OSGM-H takes $\Theta(\tfrac{1}{\eta\kappa})$ iterations to decrease the stepsize from $a$ to $\tfrac{a}{2}$ since for $\alpha_{k}\leq a$, every iteration can decrease the stepsize by at most $\eta(\kappa a-1)$:
$\alpha_{k+1}-\alpha_{k}=\eta\tfrac{(1-\alpha_{k})(x^{k}_{1})^{2}+\kappa^{2}(1-\kappa\alpha_{k})(x^{k}_{2})^{2}}{(x^{k}_{1})^{2}+\kappa^{2}(x^{k}_{2})^{2}}\geq-\eta(\kappa a-1).$
Every iteration during this phase expands $x_{2}$ by at least a factor $|1-\alpha_{k}\kappa|\geq\tfrac{a\kappa}{2}-1=\tfrac{\kappa}{6}-1$ and the spike grows as $\Omega(\kappa^{\frac{1}{\eta\kappa}})$. Hence, given any $K$, we can pick $\delta,\eta$, and $\kappa$ so that a spike appears around $K$ (Figure˜1 (right)).
2.2 Limitations of Classic-HDM
The classic version of hypergradient descent (Classic-HDM, Algorithm˜2) [2, 32] uses a different order for the primal update and stepsize update compared to OSGM-H:
$\begin{array}[]{ll}{{\texttt{Classic-HDM}}}&\texttt{OSGM-H}\\ P_{k+1}=P_{k}-\eta\nabla h_{x^{k}}(P_{k})&x^{k+1}=x^{k}-P_{k}\nabla f(x^{k})\\ x^{k+1}=x^{k}-P_{k+1}\nabla f(x^{k})\qquad&P_{k+1}=P_{k}-\eta\nabla h_{x^{k}}(P_{k})\end{array}$
To facilitate our analysis, for this section only, we drop the monotone landscape action for OSGM-H. This simplification does not compromise the intuitions we want to convey. Classic-HDM updates the stepsize using the feedback function at $x^{k}$ first and then applies the newly obtained stepsize $P_{k+1}$ to the gradient step on $x^{k}$; while OSGM-H directly applies the stepsize $P_{k}$. From an online learning perspective, the decisions in Classic-HDM utilize information from the future. Intuitively, it should yield better performance. However, the performance of Classic-HDM is inferior to OSGM-H. Figure˜2 (left) illustrates the behavior of Classic-HDM applied to a quadratic function. There are two observations:
  1. 1.
    OSGM-H converges faster than Classic-HDM under the same parameter settings.
  2. 2.
    The stepsize $\{\alpha_{k}\}$ in Classic-HDM converges, while $\{\alpha_{k}\}$ keeps oscillating in OSGM-H.
The convergence of $\{\alpha_{k}\}$ was observed by [2] and was partially explained in [5] in the case where $P$ is a full-dimensional matrix. This section further characterizes this phenomenon on a toy quadratic problem:
$\min_{x\in\mathbb{R}^{n}}\tfrac{1}{2}\langle x,Ax\rangle,$
Figure 2: Left: stepsizes $\{\alpha_{k}\}$ converge to $\alpha^{\star}=\tfrac{2}{L+\mu}$ in Classic-HDM and oscillate around $\alpha^{\star}=\tfrac{2}{L+\mu}$ in OSGM. Right: as $\alpha_{k}$ converges, Classic-HDM will converge to $x^{\star}$ along two directions.
where $A\succ 0$ has distinct eigenvalues. We will demonstrate that $\{\alpha_{k}\}$ in Classic-HDM stabilizes around the minimax optimal stepsize $\alpha^{\star}=\tfrac{2}{L+\mu}$. Besides, the iterates $\{x^{k}\}$ generated by Classic-HDM alternates between two extremal eigen directions of $A$ and lead to slow convergence. On the other hand, in OSGM-H, $\{\alpha_{k}\}$ does not stabilize and keeps adapting.
Dynamical system formulation and orbits. We can further simplify the analysis of Classic-HDM and OSGM-H by working on the basis that consists of the eigenvectors of $A$. Let $A=Q\Lambda Q^{\top}$ be the eigendecomposition. Transform the variables to $z\coloneqq Qx$. Then Classic-HDM and OSGM-H with scalar stepsize $P_{k}=\alpha_{k}I$ read as
$\begin{array}[]{ll}{{\texttt{Classic-HDM}}}&\texttt{OSGM-H}\\ \alpha_{k+1}=\alpha_{k}+\eta-\alpha_{k}\eta\tfrac{\langle z^{k},\Lambda^{3}z^{k}\rangle}{\langle z^{k},\Lambda^{2}z^{k}\rangle}\qquad&z^{k+1}=(I-\alpha_{k}\Lambda)z^{k}\\ z^{k+1}=(I-\alpha_{k+1}\Lambda)z^{k}\qquad&\alpha_{k+1}=\alpha_{k}+\eta-\alpha_{k}\eta\tfrac{\langle z^{k},\Lambda^{3}z^{k}\rangle}{\langle z^{k},\Lambda^{2}z^{k}\rangle}\end{array}$
Since the problem is homogeneous, we consider the dynamical system whose state is given by the stepsize $\alpha$ and the direction $\hat{z}\coloneqq\tfrac{z}{\|z\|}$:
[…]
Lemma 2.2 (Orbit of the dynamical system).
Suppose $A$ has distinct eigenvalues between $\mu,L]$: $L=\lambda_{1}>\lambda_{2}>\cdots>\lambda_{n-1}>\lambda_{n}=\mu$. Let $q_{i}$ denote the unit eigenvectors of $A$ associated with eigenvalues $\lambda_{i}$. Then the dynamical systems $\mathcal{F}_{{{\texttt{HDM}}}}$ and $\mathcal{F}_{\texttt{OSGM}}$ have a common period-$2$ orbit:
$\{(\alpha^{\star},\hat{z}^{\star}_{1}),(\alpha^{\star},\hat{z}^{\star}_{2})\}\coloneqq\big\{(\tfrac{2}{L+\mu},\gamma e_{1}+\delta e_{n}),(\tfrac{2}{L+\mu},-\gamma e_{1}+\delta e_{n})\big\},$
where $e_{1}=(1,0\ldots,0)$, $e_{n}=(0,\ldots,0,1)$, $\gamma=\tfrac{\mu}{\sqrt{L^{2}+\mu^{2}}}$, and $\delta=\pm\tfrac{L}{\sqrt{L^{2}+\mu^{2}}}$.
Remark 1.
Both Classic-HDM and OSGM-H have additional period-two orbits when the iterates fall in a subspace orthogonal to certain eigenvectors of $A$. For example, when $\langle q_{1},z^{1}\rangle=0$, then the orbit can be derived based on $L=\lambda_{2}$ (i.e. the maximum eigenvalue restricted on subspace). Without loss of generality, we focus on the orbit generated by the iterations that satisfy $|\langle q_{1},z^{k}\rangle|>0,|\langle q_{n},z^{k}\rangle|>0$ for all $k$. In this case, the algorithm iterations always have non-vanishing components along both top and bottom eigen directions.
Although both Classic-HDM and OSGM-H share the same orbits, Theorem˜2.1 shows that they exhibit different stability properties around the orbit. The proof of Theorem˜2.1 is deferred to appendix due to its length and technicality.
Theorem 2.1.
Let $\mathcal{J}_{{\texttt{HDM}}}$ and $\mathcal{J}_{\texttt{OSGM-H}}$ denote the Jacobians of $\mathcal{F}_{{\texttt{HDM}}}$ and $\mathcal{F}_{\texttt{OSGM}}$ respectively. For any $\eta\in(0,\frac{1}{L}]$, their spectral radius satisfy
$\rho(\mathcal{J}_{{\texttt{HDM}}}(\alpha^{\star},\hat{z}_{1}^{\star})\mathcal{J}_{{\texttt{HDM}}}(\alpha^{\star},\hat{z}_{2}^{\star}))<1\qquad\text{and}\qquad\rho(\mathcal{J}_{\texttt{OSGM}}(\alpha^{\star},\hat{z}_{1}^{\star})\mathcal{J}_{\texttt{OSGM}}(\alpha^{\star},\hat{z}_{2}^{\star}))>1$
In other words, the orbit $\{(\alpha^{\star},\hat{z}^{\star}_{1}),(\alpha^{\star},\hat{z}^{\star}_{2})\}$ is stable for $\mathcal{F}_{{{\texttt{HDM}}}}$} but unstable for $\mathcal{F}_{{{\texttt{OSGM}}}}$}.
Theorem˜2.1 indicates that $\{\alpha_{k}\}$ in Classic-HDM will get trapped around $\alpha^{\star}$ and the iterates $\{x^{k}\}$ approach the optimal solution $x^{\star}=0$ from two specific directions $d_{1}\coloneqq\gamma q_{1}+\delta q_{n}$ and $d_{2}\coloneqq-\gamma q_{1}+\delta q_{n}$ (Figure˜2 (right)). This spectral property has been extensively discussed in the literature [20, 6], and a common belief is that sticking to the orbit results in slow convergence [39]. In particular, we have
$\tfrac{\mathrm{d}}{\mathrm{d}\alpha}h_{x^{\star}+\lambda d_{1}}(\alpha^{\star})=\tfrac{\mathrm{d}}{\mathrm{d}\alpha}h_{x^{\star}+\lambda d_{2}}(\alpha^{\star})=0$
for all $\lambda\neq 0$, which means $\alpha^{\star}$ is exactly the optimal stepsize for this trajectory. As a result, Classic-HDM will stop adapting $\{\alpha_{k}\}$: Classic-HDM drives the iterates to a trajectory with no regret. However, despite Classic-HDM learns the optimal stepsize on this trajectory, this trajectory is so bad that even the optimal stepsize still offers only slow convergence. In this respect, instability of OSGM-H keeps $\{\alpha_{k}\}$ oscillating between large and small stepsizes and balances different spectral components of $x^{k}$. The idea of alternating between large and small stepsizes to leave the orbit has been explicitly used in the literature [9, 38] by designing specific stepsize schedulers. In contrast, OSGM-H automatically achieves this goal.
2.3 Heavy-ball momentum and potential function-based feedback
OSGM framework begins with the gradient-based update $x^{k+1}=x^{k}-P_{k}\nabla f(x^{k})$. In contrast, modern first-order methods with good empirical performance [23] typically add a momentum component
$x^{k+1}=x^{k}-P_{k}\nabla f(x^{k})+\beta_{k}(x^{k}-x^{k-1}),$ (5)
where $x^{k}-x^{k-1}$ is known as heavy-ball momentum [7, 29, 30] and $\beta_{k}\in\mathcal{B}\subseteq\mathbb{R}$ is the extrapolation coefficient. To adapt $(P,\beta)$ jointly, a straightforward approach defines feedback based on the objective value [2] as in OSGM-H:
$\tfrac{f(x-P\nabla f(x)+\beta(x-x^{-}))-f(x)}{\|\nabla f(x)\|^{2}}.$
However, this direct extension does not yield global convergence for two reasons: 1) due to the non-monotone behavior of standard heavy-ball momentum [7], there is no universal $(P^{\star},\beta^{\star})$ that ensures descent of the objective value; 2) a momentum method can take a large step even when the gradient norm is small, so it may happen that $\|x-x^{-}\|\gg\|\nabla f(x)\|^{2}$ and hence $\tfrac{f(x-P\nabla f(x)+\beta(x-x^{-}))-f(x)}{\|\nabla f(x)\|^{2}}$ can have an arbitrarily large Lipschitz constant $(P,\beta)$ depending on $x,x^{-}$. Feedback based on the function value alone does not faithfully reflect the quality of $(P,\beta)$ and cannot guide the scheduler constructively. To identify a suitable feedback, recall that heavy-ball momentum admits a potential function [7]
$\varphi_{\omega}(x,x^{-})\coloneqq f(x)-f^{\star}+\tfrac{\omega}{2}\|x-x^{-}\|^{2}.$ (6)
This potential function monotonically decreases if $(P,\beta)$ are appropriately chosen.
Lemma 2.3 (Heavy-ball potential [7, Theorem 2]).
Suppose $f$ is $L$-smooth and convex. Given any $x,x^{-}\not\in\mathcal{X}^{\star}$, denote $x^{+}\coloneqq x-P\nabla f(x)+\beta(x-x^{-})$ with $P=\alpha I$, $\alpha\leq\tfrac{1}{L}$, and $\beta\in[0,\sqrt{1-\alpha L}]$. Then the potential $\varphi_{\omega}$ with $\omega=\tfrac{1-\alpha L}{2\alpha}$ satisfies
$\varphi_{\omega}(x^{+},x)\leq\varphi_{\omega}(x,x^{-})-\tfrac{\alpha}{2}\big[\|\nabla f(x)\|^{2}+\tfrac{1-\alpha L-\beta^{2}}{\alpha^{2}}\|x-x^{-}\|^{2}\big].$
Lemma˜2.3 resembles the descent lemma: the progress of each heavy-ball step in reducing $\varphi_{\omega}$ is proportional to $\|\nabla f(x)\|^{2}+\tfrac{\tau}{2}\|x-x^{-}\|^{2}$ for $\tau=\tfrac{1-\alpha L-\beta^{2}}{\alpha^{2}}>0$. Therefore, we define the heavy-ball feedback as
$h_{x,x^{-}}(P,\beta)\coloneqq\tfrac{\varphi_{\omega}(x^{+}(P,\beta),x)-\varphi_{\omega}(x,x^{-})}{\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}},$ (7)
where $x^{+}(P,\beta)=x-P\nabla f(x)+\beta(x-x^{-})$ and $\tau>0$ is a parameter.
The heavy-ball feedback (7) depends on two points $x$ and $x^{-}$, which we denote by a state $z=(z_{1},z_{2})=(x,x^{-})$. With slight abuse of notation, we denote $f(z):=f(z_{1})$ and $\nabla f(z):=\nabla f(z_{1})$. We will also slightly abuse the notation of the gradient by directly concatenating the gradients with respect to $P$ and $\beta$: $\nabla h_{z}(P,\beta)=(\nabla_{P}h_{z}(P,\beta),\nabla_{\beta}h_{z}(P,\beta))$. The pushforward operation and the heavy-ball feedback are written with respect to state $z$ as
$z^{+}(P,\beta)\coloneqq(z_{1}-P\nabla f(z)+\beta(z_{1}-z_{2}),~z_{1}),\qquad h_{z}(P,\beta)\coloneqq\tfrac{\varphi_{\omega}(z^{+}(P,\beta))-\varphi_{\omega}(z)}{\|\nabla f(z)\|^{2}+\frac{\tau}{2}\|z_{1}-z_{2}\|^{2}}.$
Using the superscript $k$ for the iteration count, the first component of the state $z^{k}\coloneqq(z_{1}^{k},z_{2}^{k})$ satisfies $z^{k}_{1}=x^{k}$; however, the second component $z^{k}_{2}$ need not be equal to $x^{k-1}$. We will utilize this flexibility to implement the monotone landscape action outlined below. In the product space $\mathcal{P}\times\mathcal{B}$ of stepsize and momentum, define the norm and its associated dual norm as
$\|(P,\beta)\|\coloneqq\sqrt{\|P\|_{F}^{2}+\tfrac{1}{L^{2}}\beta^{2}},\qquad\|(P,\beta)\|_{\ast}\coloneqq\sqrt{\|P\|_{F}^{2}+L^{2}\beta^{2}}.$ (8)
The weight $\tfrac{1}{L^{2}}$ ensures both $\|P\|_{F}^{2}$ and $\tfrac{1}{L^{2}}\beta^{2}$ have the same physical dimension since $P$ lies in the space of inverse Hessian and $\beta$ is dimensionless. We prove properties of feedback $h_{z}(P,\beta)$ with respect to this norm.
Lemma 2.4 (Properties of heavy-ball feedback).
Suppose $f$ is convex and $L$-smooth and let $\tau\geq 2L^{2}$. Then for any state $z=(z_{1},z_{2})$ with $z_{1},z_{2}\not\in\mathcal{X}^{\star}$, the heavy-ball feedback $h_{z}(P,\beta)$ is jointly convex in $(P,\beta)$ and $(L+\omega)$-smooth in $(P,\beta)$. Moreover, the gradient of $h_{z}(P,\beta)$ takes the form
$\nabla h_{z}(P,\beta)=\Big(-\tfrac{[\nabla f(z^{+}(P,\beta))+\omega(-P\nabla f(z)+\beta(z_{1}-z_{2}))]\nabla f(z)^{\top}}{\|\nabla f(z)\|^{2}+\frac{\tau}{2}\|z_{1}-z_{2}\|^{2}},~\tfrac{\langle[\nabla f(z^{+}(P,\beta))+\omega(-P\nabla f(z)+\beta(z_{1}-z_{2}))],~z_{1}-z_{2}\rangle}{\|\nabla f(z)\|^{2}+\frac{\tau}{2}\|z_{1}-z_{2}\|^{2}}\Big).$
At each iteration, define the progress with respect to the heavy-ball feedback:
$b_{k}\coloneqq\tfrac{\varphi_{\omega}(z^{k+1})-\varphi_{\omega}(z^{k})}{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}.$
Theorem˜2.2 shows a reduction from the function value gap to the cumulative progress $\sum^{K}_{k=1}b_{k}$.
Theorem 2.2 (Heavy-ball reduction).
Choose the parameters $\omega>0$ and $\tau>0$ such that $\tau\geq 2L\omega$. Let $\{z^{k}\}$ be any sequence of states such that $\varphi_{\omega}(z^{k+1})\leq\varphi_{\omega}(z^{k})$ and let the initial state $z^{1}=(z_{1}^{1},z_{2}^{1})$ satisfy $z_{1}^{1}=z_{2}^{1}$.
  • If $f$ is convex, then
    $f(z^{K+1})-f^{\star}\leq\tfrac{f(z^{1})-f^{\star}}{1+V\sum^{K}_{k=1}-b_{k}},$ (9)
    where $V\coloneqq\min\big\{\tfrac{f(z^{1})-f^{\star}}{4\Delta^{2}},\tfrac{L}{2}\big\}$ and $\Delta\coloneqq\max_{x\in\{x:f(x)\leq f(x^{1})\}}\min_{x^{\star}\in\mathcal{X}^{\star}}\|x-x^{\star}\|$.
  • If $f$ is $\mu$-strongly convex, then
    $f(z^{K+1})-f^{\star}\leq[f(z^{1})-f^{\star}](1-\tfrac{\mu}{K}\sum_{k=1}^{K}-b_{k})^{K}.$ (10)
The heavy-ball reduction (Theorem˜2.2) requires monotonicity of the potential $\varphi_{\omega}(z^{k})$, instead of function value $f(z^{k})$. Hence, given the proposal state from the scheduler
$z^{k+1/2}\coloneqq(z^{k})^{+}=(z_{1}^{k}-P_{k}\nabla f(z^{k})+\beta_{k}(z_{1}^{k}-z_{2}^{k}),z_{1}^{k}),$ (11)
the monotone landscape action for heavy-ball feedback must ensure the potential is nonincreasing:
$\varphi_{\omega}(z^{k+1})\leq\min\{\varphi_{\omega}(z^{k+1/2}),\varphi_{\omega}(z^{k})\},$
where $z^{k+1}$ is the state returned by the monotone landscape action. We may implement this monotone landscape action by the null step $z^{k+1}=\operatornamewithlimits{arg\,min}\big\{\varphi_{\omega}(z^{k+1/2}),\varphi_{\omega}(z^{k})\big\}$, which simply returns whichever state has the smaller potential value. In practice, the lookahead action described in Section˜2.4 is often preferred.
2.4 Algorithm design of OSGM-Best
We put together the insights above and propose OSGM-Best, a variant of OSGM that demonstrates competitive performance in practice. OSGM-Best utilizes the lookahead landscape action, tailored to heavy-ball feedback, to suppress regret and thereby accelerate convergence.
Part I [14] introduced the lookahead landscape action, which takes an additional gradient step on the proposed iterate $x^{k+1/2}$ to suppress the regret of online gradient descent on feedback. Since the heavy-ball feedback is defined with respect to the potential function, the additional gradient step should be taken with respect to the potential function $\varphi_{\omega}(\cdot,~z_{2}^{k+1/2})$. As a result, the lookahead landscape action for heavy-ball feedback is
$\mathcal{M}_{\text{look}}(z^{k+1/2})$ $\coloneqq{}(z_{1}^{k+1/2}-\tfrac{1}{L+\omega}\nabla_{1}\varphi_{\omega}(z^{k+1/2}),~z_{2}^{k+1/2})$
$=(z_{1}^{k+1/2}-\tfrac{1}{L+\omega}[\nabla f(z^{k+1/2})+\omega(z_{1}^{k+1/2}-z_{2}^{k+1/2})],~z_{2}^{k+1/2})$
$=((1-\tfrac{\omega}{L+\omega})z_{1}^{k+1/2}+\tfrac{\omega}{L+\omega}z_{2}^{k+1/2}-\tfrac{1}{L+\omega}\nabla f(z^{k+1/2}),~z_{2}^{k+1/2}),$
where $\nabla_{1}\varphi_{\omega}(z)$ denotes the gradient with respect to $z_{1}$ and the stepsize $\frac{1}{L+\omega}$ is the inverse of the smoothness constant $L+\omega$ of the potential function $\varphi_{\omega}(z)$ with respect to $z_{1}$. OSGM-Best determines the next state $z^{k+1}$ by composing the lookahead and monotone landscape actions:
$\varphi_{\omega}(z^{k+1})\leq\min\{\varphi_{\omega}(\mathcal{M}_{\text{look}}(z^{k+1/2})),\varphi_{\omega}(z^{k})\},$ (12)
Algorithm˜3 summarizes OSGM-Best with monotone landscape action implemented by the null step. We refer to Algorithm˜3 as OSGM-Best since it is empirically superior to other OSGM variants. The global convergence of OSGM is established in Theorem˜2.3.
1input Initial state $z^{1}=(z^{1}_{1},z^{1}_{2})$ with $z_{1}^{1}=z_{2}^{1}$, initial stepsize $P_{1}\in\mathbb{R}^{n\times n}$ and momentum parameter $\beta_{1}\in\mathbb{R}$, online gradient descent stepsizes $\eta_{P},\eta_{\beta}>0$
2for $k=1,2,\dots$ do
3 $z^{k+1/2}=(z_{1}^{k}-P_{k}\nabla f(z^{k})+\beta_{k}(z_{1}^{k}-z_{2}^{k}),~z_{1}^{k})$ $\triangleright$  State proposal from scheduler
4 $z_{\text{look}}^{k+1}=(z_{1}^{k+1/2}-\tfrac{1}{L+\omega}\nabla_{1}\varphi_{\omega}(z^{k+1/2}),~z_{2}^{k+1/2})$ $\triangleright$  Lookahead action
5   Choose $z^{k+1}$ s.t. $\varphi_{\omega}(z^{k+1})\leq\min\{\varphi_{\omega}(z_{\text{look}}^{k+1}),\varphi_{\omega}(z^{k})\}$ $\triangleright$  Monotone action by null step
6 $P_{k+1}=P_{k}-\eta_{P}\nabla_{P}h_{z^{k}}(P_{k},\beta_{k})$ $\triangleright$  Online gradient update on $P$
7 $\beta_{k+1}=\beta_{k}-\eta_{\beta}\nabla_{\beta}h_{z^{k}}(P_{k},\beta_{k})$ $\triangleright$  Online gradient update on $\beta$
8
9 end for
Algorithm 3 OSGM-Best
Theorem 2.3 (Global convergence).
Suppose $f$ is $L$-smooth. Choose $\omega=3L$, $\tau=16L^{2}$ and online gradient stepsizes $\eta_{P}=\tfrac{\eta_{\beta}}{L^{2}}=\frac{1}{2L}$. Then for any benchmark stepsize $\hat{P}\in\mathbb{R}^{n\times n}$ and momentum parameter $\hat{\beta}\in\mathbb{R}$,
$f(x^{K+1})-f^{\star}$ $\leq\frac{f(x^{1})-f^{\star}}{KV\max\{\frac{1}{K}\textstyle\sum_{k=1}^{K}-h_{z^{k}}(\hat{P},\hat{\beta})-\frac{L}{4K}\|(P_{1}-\hat{P},\beta_{1}-\hat{\beta})\|^{2},0\}+1},$ (convex)
$f(x^{K+1})-f^{\star}$ $\leq[f(x^{1})-f^{\star}](1-\mu\max\{\tfrac{1}{K}\textstyle\sum_{k=1}^{K}-h_{z^{k}}(\hat{P},\hat{\beta})-\tfrac{L}{4K}\|(P_{1}-\hat{P},\beta_{1}-\hat{\beta})\|^{2},0\})^{K}.$ ($\mu$-strongly convex)
where $V$ and $\Delta$ are constants defined in Theorem˜2.2. In particular, with $(P_{1},\beta_{1})=(\frac{1}{4L}I,\frac{1}{2})$, we have
$f(x^{K+1})-f^{\star}$ $\leq\tfrac{8L[f(x^{1})-f^{\star}]}{KV},$ (convex)
$f(x^{K+1})-f^{\star}$ $\leq[f(x^{1})-f^{\star}](1-\tfrac{1}{8\kappa})^{K}.$ ($\mu$-strongly convex)
3 OSGM for smooth nonconvex optimization
OSGM can be extended to smooth nonconvex optimization problems, but requires modifying the feedback. The analysis of OSGM in Part I [14] and Section˜2 assumes a convex objective $f$, which guarantees convexity of feedback and enables hypergradient [14, Theorem 4.2] and heavy-ball (Theorem˜2.2) progress reduction. One possible approach to handle nonconvex $f$ is to apply the proximal point method [4] with a sufficiently large regularizer $\gamma>L$:
$x^{k+1}=\operatornamewithlimits{arg\,min}_{x}~\{f(x)+\tfrac{\gamma}{2}\|x-x^{k}\|^{2}\}.$ (13)
When $f$ is nonconvex but $L$-smooth, the proximal subproblem (13) is $(L-\gamma)$-weakly convex and becomes strongly convex if $\gamma>L$. Then OSGM can serve as a subroutine to solve the subproblem (13). However, this proximal point approach requires a double-loop algorithm: an outer loop updates the proximal center $x^{k}$ and an inner loop solves the proximal subproblem with OSGM. In contrast, this section introduces a different approach to smooth nonconvex optimization by stepsize regularization, avoiding the need for a double-loop algorithm. Observe that $\tfrac{1}{L}I$ is a safe stepsize in smooth nonconvex optimization. Hence, we can regularize the feedback function $h_{x}(P)$ by $\|P-\tfrac{1}{L}I\|_{F}^{2}$. This regularizer vanishes when the stepsize $\hat{P}=\frac{1}{L}I$ is used as a benchmark, and an online learning method using this regularized feedback at least competes with the performance of $\tfrac{1}{L}I$. Regularization in stepsize space requires only a fixed center $\tfrac{1}{L}I$ and avoids the second loop.
Structure of the section.
The roadmap of this section resembles Part I: Section˜3.1 introduces the regularized feedback and establishes its analytical properties; Section˜3.2 introduces two progress reductions for general nonconvex and PL functions and several landscape actions. Finally, in Section˜3.3 we design and analyze OSGM-H variants for nonconvex optimization. For simplicity, we will again focus on the hypergradient feedback, while the results for ratio feedback hold similarly for nonconvex PL functions.
3.1 Regularized feedback design and analysis
Given a regularization parameter $\lambda\geq 0$, define the regularized hypergradient feedback
$h_{x}^{\lambda}(P)\coloneqq h_{x}(P)+\tfrac{\lambda}{2}\|P-\tfrac{1}{L}I\|_{F}^{2}=\tfrac{f(x-P\nabla f(x))-f(x)}{\|\nabla f(x)\|^{2}}+\tfrac{\lambda}{2}\|P-\tfrac{1}{L}\|_{F}^{2}.$ (14)
The regularization parameter $\lambda$ is chosen to neutralize the local nonconvexity of $h_{x}$ for $P\in\mathcal{P}$. Since $h_{x}(P)$ is $L$-smooth, choosing $\lambda=L$ suffices since it guarantees $h^{\lambda}_{x}$ is globally convex. However, this strong regularizer forces $P$ to stay close to the conservative stepsize $\tfrac{1}{L}I$. In practice, we want to choose $\lambda$ as small as possible to retain adaptivity, especially when $x$ enters a locally convex region. In other words, we need an estimate for the local weak convexity constant of $h_{x}(P)$. With additional assumptions on $f$, it is possible.
Lemma 3.1 (Nonconvexity of $h_{x}$).
For any $x\not\in\mathcal{S}^{\star}\coloneqq\{x:\|\nabla f(x)\|=0\}$, $h_{x}(P)$ has the following properties:
  1. 1.
    If $f$ is $L$-smooth, then $h_{x}(P)$ has weak convexity constant $\rho\leq\rho_{1}\coloneqq L$.
  2. 2.
    If $f$ is $L$-smooth, has $H$-Lipschitz Hessian and $\mathrm{diam}(\mathcal{P})\leq D$, then $h_{x}(P)$ has weak convexity constant
    $\rho\leq\rho_{2}\coloneqq\max\{-\lambda_{\min}(\nabla^{2}h_{x}(0\cdot I))+HD\|\nabla f(x)\|,0\}$
    when $P\in\mathcal{P}$.
  3. 3.
    Moreover, if $f$ is also $\mu$-PL, then $\rho\leq\rho_{3}\coloneqq(\tfrac{H}{\mu}+HD)\|\nabla f(x)\|$.
The intuition behind Lemma˜3.1 is simple: when $x$ is close to stationarity and $\mathrm{diam}(\mathcal{P})<\infty$, then $\nabla^{2}h_{x}(P)$ does not change much within $P\in\mathcal{P}$. Hence, we can bound the worst-case weak convexity constant of $h_{x}(P)$ for any $P\in\mathcal{P}$ by evaluating $\lambda_{\min}(\nabla^{2}h_{x}(\bar{P}))$ for some fixed $\bar{P}\in\mathcal{P}$. In particular, if $f$ is $\mu$-PL, $\lambda_{\min}(\nabla^{2}h_{x}(\bar{P}))$ can also be bounded using $\|\nabla f(x)\|$. Lemma˜3.1 takes $\bar{P}=0\cdot I$.
Remark 2.
Computing $\lambda_{\min}(\nabla^{2}h_{x}(\bar{P}))$ can be computationally expensive in practice. One exception is when $P=\alpha I$ is a scalar multiple of the identity. In this case the second-order derivative of $h_{x}(\alpha I)$ with respect to $\alpha$ is $\langle\tfrac{\nabla f(x)}{\|\nabla f(x)\|},\nabla^{2}f(x-\alpha\nabla f(x))\tfrac{\nabla f(x)}{\|\nabla f(x)\|}\rangle$ and can be computed by evaluating the Hessian-vector product, without forming the Hessian matrix [28].
Since $h_{x}^{\lambda}(P)$ adds a strongly convex regularizer to $h_{x}(P)$, its analytic properties follow from those of $h_{x}(P)$.
Lemma 3.2 (Properties of $h_{x}^{\lambda}$).
For any $x\not\in\mathcal{S}^{\star}$, the regularized feedback $h_{x}^{\lambda}(P)$ has gradient
$\nabla h_{x}^{\lambda}(P)=-\tfrac{\nabla f(x-P\nabla f(x))\nabla f(x)^{\top}}{\|\nabla f(x)\|^{2}}+\lambda(P-\tfrac{1}{L}I).$
and $h_{x}^{\lambda}(\tfrac{1}{L}I)\leq-\tfrac{1}{2L}$ for all $\lambda$. Moreover, if $f$ is $L$-smooth, has $H$-Lipschitz Hessian and $\mathrm{diam}(\mathcal{P})\leq D$, then $h_{x}^{\lambda}(P)$ is convex and $(L+\lambda)$-smooth for all $\lambda\geq\min\{\rho_{1},\rho_{2},\rho_{3}\}$.
Finally, the definition of minimax stepsize does not immediately carry over to the nonconvex case; therefore, we use $\frac{1}{L}I$ as the benchmark stepsize in the theoretical analyses.
3.2 Nonconvex progress reduction and landscape action
Recall the per iteration progress associated with the hypergradient feedback $h_{k}\coloneqq\tfrac{f(x^{k+1})-f(x^{k})}{\|\nabla f(x^{k})\|^{2}}$. Theorem˜3.1 reduces the stationarity/optimality guarantees to the cumulative progress. Notably, these two reductions do not assume the convexity of $f$.
Theorem 3.1 (Nonconvex reductions).
Let $\{x^{k}\}$ be any sequence such that $x^{k}\not\in\mathcal{S}^{\star}$ and $f(x^{k+1})\leq f(x^{k})$. Then
$\min_{1\leq k\leq K}\|\nabla f(x^{k})\|^{2}\leq\tfrac{f(x^{1})-f^{\star}}{K}\tfrac{1}{\frac{1}{K}\sum_{k=1}^{K}-h_{k}}.$ (15)
Moreover, if $f$ is $\mu$-PL, then
$f(x^{K+1})-f^{\star}\leq[f(x^{1})-f^{\star}](1-\frac{2\mu}{K}\sum_{k=1}^{K}-h_{k})^{K}.$ (16)
Landscape actions $\mathcal{M}$ and progress.
As in Part I, the landscape action $\mathcal{M}(x^{k+1/2},x^{k})$ determines the choice of the next iterate in OSGM. Recall the two landscape actions in OSGM-H:
  • Monotone. $x^{k+1}$ satisfies $f(x^{k+1})\leq\min\{f(x^{k+1/2}),f(x^{k})\}$.
  • Monotone lookahead. $x^{k+1}$ satisfies $f(x^{k+1})\leq\min\{f(x^{k+1/2}-\tfrac{1}{L}\nabla f(x^{k+1/2})),f(x^{k})\}$.
For each action, we associate the progress and feedback in Lemma˜3.3.
Lemma 3.3 (Feedback and progress).
Let $f$ be $L$-smooth and suppose $\lambda_{k}\leq L$. Then each of the above two landscape actions guarantees the following relation between the feedback and per iteration progress:
  • Monotone. $h_{k}\leq\min\{h^{\lambda_{k}}_{x^{k}}(P_{k})-\tfrac{\lambda_{k}}{2}\|P_{k}-\tfrac{1}{L}\|_{F}^{2},0\}$.
  • Monotone lookahead. $h_{k}\leq\min\{h^{\lambda_{k}}_{x^{k}}(P_{k})-\tfrac{1}{4L}\|\nabla h^{\lambda_{k}}_{x^{k}}(P_{k})\|_{F}^{2},0\}$.
3.3 Algorithm design and analysis
Putting the results together, we instantiate Monotone Lookahead OSGM-H for nonconvex optimization with
$\ell_{x}(P)\coloneqq h_{x}^{\lambda}(P),\quad\text{Monotone Lookahead landscape: }f(x^{k+1})\leq\min\{f(x^{k+1/2}-\tfrac{1}{L}\nabla f(x^{k+1/2})),f(x^{k})\}\quad\mathcal{A}\coloneqq{\texttt{OGD}}.$
The algorithm is Monotone Lookahead OSGM-H with regularization (Algorithm˜4).
1input initial point $x^{1}$, initial stepsize $P_{1}\in\mathcal{P}$, online gradient stepsize $\eta>0$
2for $k=1,2,\dots$ do
3 $x^{k+1/2}=x^{k}-P_{k}\nabla f(x^{k})$
4   Choose $x^{k+1}$ that satisfies $f(x^{k+1})\leq\min\{f(x^{k+1/2}-\frac{1}{L}\nabla f(x^{k+1/2})),f(x^{k})\}$
5   Choose $0\leq\lambda_{k}\leq L$ based on Lemma˜3.1 such that $h_{x^{k}}^{\lambda_{k}}(P)$ is convex within $\mathcal{P}$
6 $P_{k+1}=\Pi_{\mathcal{P}}[P_{k}-\eta\nabla h_{x^{k}}^{\lambda_{k}}(P_{k})]$
7 end for
Algorithm 4 Monotone Lookahead OSGM-H for nonconvex optimization
Recall that the gradient of the regularized hypergradient feedback in Algorithm˜4 of Algorithm˜4 takes the form
$\nabla h_{x}^{\lambda}(P)=-\tfrac{\nabla f(x-P\nabla f(x))\nabla f(x)^{\top}}{\|\nabla f(x)\|^{2}}+\lambda(P-\tfrac{1}{L}I)$
and $\mathcal{P}$ can be chosen as a simple bounded set (e.g., an element-wise box) to allow inexpensive orthogonal projection. Due to the difficulty of the nonconvex setting, we restrict our analysis to global convergence only.
Theorem 3.2 (Convergence).
Let $f$ be $L$-smooth. For any benchmark stepsize $\hat{P}\in\mathcal{P}$, Monotone Lookahead OSGM-H (Algorithm˜4) with $\eta=1/(4L)$ satisfies
$\min_{1\leq k\leq K}\|\nabla f(x^{k})\|^{2}\leq{}$ $\tfrac{f(x^{1})-f^{\star}}{K}\tfrac{1}{\frac{1}{K}\sum_{k=1}^{K}[-h_{x^{k}}(\hat{P})-\frac{\lambda_{k}}{2}\|\hat{P}-\frac{1}{L}I\|_{F}^{2}]-\frac{L}{K}\|P_{1}-\hat{P}\|_{F}^{2}}$ (nonconvex)
$f(x^{K+1})-f^{\star}\leq{}$ $\textstyle[f(x^{1})-f^{\star}](1-2\mu\max\{\tfrac{1}{K}\sum_{k=1}^{K}[-h_{x^{k}}(\hat{P})-\tfrac{\lambda_{k}}{2}\|\hat{P}-\tfrac{1}{L}I\|_{F}^{2}]-\tfrac{L}{K}\|P_{1}-\hat{P}\|_{F}^{2},0\})^{K}.$
In particular, if $P_{1}=\tfrac{1}{L}I$, then
$\min_{1\leq k\leq K}\|\nabla f(x^{k})\|^{2}\leq{}$ $\tfrac{2L[f(x^{1})-f^{\star}]}{K}$ (nonconvex)
$f(x^{K+1})-f^{\star}\leq{}$ $[f(x^{1})-f^{\star}](1-\tfrac{1}{\kappa})^{K}.$
Theorem˜3.2 does not show an improved convergence rate compared to gradient descent using constant stepsize $\frac{1}{L}I$. However, when the iterates enter a locally convex region (e.g., approaching a local minimum), $\lambda_{k}$ will automatically vanish, and we recover adaptivity in the convex case.
4 Numerical experiments
This section conducts numerical experiments to validate the empirical performance of hypergradient descent. We compare OSGM-Best with different adaptive optimization algorithms.
4.1 Experiment setup
Dataset.
We use 38 datasets from LIBSVM for convex optimization tasks and 47 unconstrained problems from CUTEst dataset [16].
Algorithm benchmark.
We benchmark the following algorithms.
  • GD. Vanilla gradient descent.
  • GD-HB. Gradient descent with heavy-ball momentum. [29]
  • AGD-CVX. The smooth convex version of accelerated gradient descent (Nesterov momentum). [8]
  • AGD-SCVX. The smooth strongly convex version of accelerated gradient descent. [8]
  • Adam. Adaptive momentum estimation. [23]
  • AdaGrad. Adaptive (sub)gradient method. [10]
  • BFGS. BFGS from scipy [26, 36].
  • L-BFGS-Mk. L-BFGS with memory size k in scipy.
  • Practical variant OSGM-Best uses as memory $7$ vectors of size $n$, comparable to memory for L-BFGS-M1.
Algorithm configuration.
  • Stepsize in GD, GD-HB, AGD-CVX, and AGD-SCVX are all set to $1/L$.
  • The momentum parameter in GD-HB is chosen within the set $\{0.1,0.5,0.9,0.99\}$.
  • The Adam stepsize is chosen within the set $\{1/L,10^{-3},10^{-2},10^{-1},1,10\}$. $\beta_{1}=0.9,\beta_{2}=0.999$.
  • The AdaGrad stepsize is chosen within the set $\{1/L,10^{-3},10^{-2},10^{-1},1,10\}$.
  • BFGS, L-BFGS-Mk use default parameters in scipy.
  • OSGM-Best uses the default parameters without tuning.
Testing configurations.
  1. 1)
    Maximum oracle access. For convex problems, we allow a maximum of 1000 gradient oracles for each algorithm. For nonconvex problems, we allow a maximum of 2000 gradient oracles.
  2. 2)
    Initial point. For convex problems, All the algorithms are initialized from the same starting point generated from the normal distribution $\mathcal{N}(0,I_{n})$ and normalized to have unit length. For nonconvex problems, we use the starting point provided by CUTEst.
  3. 3)
    Stopping criterion. Algorithms stop if $\|\nabla f\|_{\infty}\leq 10^{-3}$.
Table 1: Number of solved problems for each algorithm.
Algorithm/Problem SVM (38) $\uparrow$ Logistic Regression (38) $\uparrow$ CUTEst (47) $\uparrow$
GD 5 2 13
GD-HB 13 10 20
AGD-CVX 13 6 20
AGD-SCVX 10 9 15
Adam 31 16 24
AdaGrad 11 11 18
L-BFGS-M1 27 18 35
L-BFGS-M3 33 19 37
L-BFGS-M5 35 26 37
L-BFGS-M10 35 31 36
BFGS 37 36 38
OSGM-Best 35 35 34
4.2 Convex optimization
Figure 3: Support vector-machine problems. First row: function value gap. Second row: gradient norm.
Figure 4: Logistic regression problems. First row: function value gap. Second row: gradient norm.
For each algorithm, we record the number of successfully solved instances ($\|\nabla f\|_{\infty}\leq 10^{-3}$ within 1000 gradient oracles). Table˜1 summarizes the detailed statistics. The number of instances solved by OSGM-Best is comparable to that of L-BFGS-M10.
Support Vector Machine.
Figure˜5 shows the function value gap and gradient norm plots on sample test instances on support vector machine problems. The optimal value for each instance is obtained by running BFGS until $\|\nabla f\|_{\infty}\leq 10^{-3}$. We see that the practical variant of OSGM-Best achieves a significant speedup over other adaptive first-order methods. In particular, OSGM-Best often matches L-BFGS-M5 and L-BFGS-M10, while its memory usage is closer to L-BFGS-M1. Notably, Adam also achieves competitive performance in several instances.
Logistic Regression.
In logistic regression (Figure˜4), OSGM-Best still compares well with L-BFGS-M5 and is significantly faster than other adaptive first-order methods.
4.3 Nonconvex optimization
On nonconvex problems, OSGM-Best still significantly outperforms standard adaptive first-order methods, and its performance is comparable to L-BFGS.
Figure 5: CUTEst problems. First row: function value gap. Second row: gradient norm.
5 Future directions and open problems
In this section, we sketch extensions of OSGM and provide several possible future directions.
5.1 Barzilai-Borwein stepsize and online proximal point
The Barzilai-Borwein stepsize (also known as the BB step) is a widely used practical stepsize selection strategy [1]. There are two types of BB steps:
$\alpha_{k}\coloneqq\tfrac{\|x^{k}-x^{k-1}\|^{2}}{\langle\nabla f(x^{k})-\nabla f(x^{k-1}),x^{k}-x^{k-1}\rangle}\quad\text{and}\quad\alpha^{\prime}_{k}\coloneqq\tfrac{\langle\nabla f(x^{k})-\nabla f(x^{k-1}),x^{k}-x^{k-1}\rangle}{\|\nabla f(x^{k})-\nabla f(x^{k-1})\|^{2}}.$
In the context of convex quadratic minimization, it is known that $\alpha_{k}$ is the steepest descent stepsize at the previous iteration in terms of function value [39, Section 2.1],
$\alpha_{k}=\operatornamewithlimits{arg\,min}_{\alpha}f(x^{k-1}-\alpha\nabla f(x^{k-1})),$
and similarly, $\alpha^{\prime}_{k}$ is the steepest descent stepsize at the previous iteration in terms of the gradient norm
$\alpha_{k}^{\prime}=\operatornamewithlimits{arg\,min}_{\alpha}\|\nabla f(x^{k-1}-\alpha\nabla f(x^{k-1}))\|.$
Both BB steps fit into OSGM framework by rewriting them as the minimizers of feedback functions:
$\alpha_{k+1}=\operatornamewithlimits{arg\,min}_{\alpha}~h_{x^{k}}(\alpha)\coloneqq\tfrac{f(x^{k}-\alpha\nabla f(x^{k}))-f(x^{k})}{\|\nabla f(x^{k})\|^{2}}\quad\text{and}\quad\alpha^{\prime}_{k+1}=\operatornamewithlimits{arg\,min}_{\alpha}~g_{x^{k}}(\alpha)\coloneqq\tfrac{\|\nabla f(x^{k}-\alpha\nabla f(x^{k}))\|}{\|\nabla f(x^{k})\|},$
where $h_{x}(\alpha)$ is the hypergradient feedback and $g_{x}(\alpha)$ is the gradient norm feedback discussed in [13]. The BB steps can be derived from the minimization of one of the feedback functions, and we may further develop this perspective to create variants of the BB step powered by online learning. For example, we can write the BB step as as the online proximal point update (also known as implicit online learning [24]) with $\eta\rightarrow\infty$
$\alpha_{k+1}=\operatornamewithlimits{arg\,min}_{\alpha}~\{h_{x^{k}}(\alpha)+\tfrac{1}{2\eta}(\alpha-\alpha_{k})^{2}\}.$ (17)
To our knowledge, the connection between the BB step and online learning has not been previously explored. Related ideas have primarily appeared in the literature under the guise of “retarded” or “delayed” gradient updates [12, 39]. We believe this new perspective can motivate extensions or improvements on the BB step. For example, using finite values of $\eta$ above would yield stabilized algorithms similar to the BB step, with a pathway towards a provable convergence guarantee through this connection to the online proximal point method and OSGM. Alternatively, this perspective clarifies how to develop momentum or diagonal variants of the BB step. In Section˜C.1, we illustrate a convergence analysis of OSGM-H based on (17).
Question 1.
The BB step is guaranteed to converge on quadratics, but no theoretical guarantees are available for general functions without complex linesearch [37]. Is it possible to use this online learning perspective to design a competitive variant of the BB step with a theoretically guaranteed convergence rate for general functions?
Finally, for quadratic $f$, the BB step uses the steepest descent stepsize from the previous iteration. In contrast, the steepest descent method knows the future ahead of time and applies the “optimal” stepsize. Although the BB step can be unstable when $f$ has a large condition number, it typically outperforms steepest descent. This phenomenon, where knowing the future hinders convergence, aligns with our observation in Section˜2.2. To formally state the analogy, steepest descent is to BB step what Classic-HDM is to OSGM-H.
5.2 OSGM-R without knowing $f^{\star}$
Although OSGM-H has more appealing practical performance, only OSGM-R theoretically improves the complexity of gradient descent. However, OSGM-R has a drawback of requiring knowledge of $f^{\star}$.
Question 2.
Is it possible to obtain $\mathcal{O}(\kappa^{\star}\log(1/\varepsilon))$ complexity without knowledge of $f^{\star}$?
This issue also appears in well-known methods such as the Polyak stepsize [19]. Two strategies from the Polyak stepsize literature can be borrowed: a double-loop algorithm that searches for $f^{\star}$ can maintain a provable convergence rate, but introduces an additional $\log(1/\varepsilon)$ factor in the bound [19]. On the other hand, methods that dynamically estimate $f^{\star}$, such as the variable target or moving target methods [17, 22], typically cannot recover a nonasymptotic convergence rate in the smooth strongly convex case. Here, we demonstrate an additional (partial) solution to this question in the special case when a dual bound on the primal objective $f$ can be obtained using the Fenchel trick. Suppose the original optimization problem takes the form
$\min_{x\in\mathbb{R}^{n}}~f(x)\coloneqq f_{1}(x)+f_{2}(Ax)$
where $A\in\mathbb{R}^{m\times n}$. Then, using the Fenchel dual trick, we can write the dual problem as
$\max_{y\in\mathbb{R}^{m}}~g(y)\coloneqq-f_{1}^{\ast}(-A^{\top}y)-f_{2}^{\ast}(y).$
Suppose $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and $g:\mathbb{R}^{m}\rightarrow\mathbb{R}$ are both smooth strongly convex, and that first-order oracles $f(x),\nabla f(x),g(y),\nabla g(y)$ can be efficiently evaluated. Examples satisfying these two conditions include ridge regression, where $f(x)=\tfrac{1}{2}\|Ax-b\|^{2}+\tfrac{\lambda}{2}\|x\|^{2},g(y)=-\tfrac{1}{2\lambda}\|A^{\top}y\|^{2}-\langle b,y\rangle-\tfrac{1}{2}\|y\|^{2}$ and certain regularized generalized linear models [33]. Then strong duality holds $\min_{x}f(x)=f^{\star}=\max_{y}g(y)$, and we define the duality gap function $F(z)=F(x,y)\coloneqq f(x)-g(y)$. Since $F(z^{\star})=0$, we can apply OSGM to the duality gap function $F$ to achieve $\mathcal{O}(\kappa^{\star}_{F}\log(1/\varepsilon))$ complexity, where $\kappa^{\star}_{F}$ is the optimal condition number of the gap function. This idea can also be applied to remove $f^{\star}$ in the Polyak stepsize.
5.3 OSGM for accelerated gradient descent
OSGM works well with heavy-ball momentum. Can it similarly improve the performance of Nesterov momentum, for example, by using AGD as our base algorithm instead of gradient descent?
Question 3.
Is there a method that (with knowledge of $f^{\star}$) achieves (asymptotic) $\mathcal{O}(\sqrt{\kappa^{\star}}\log(1/\varepsilon))$ complexity with only access to a first-order oracle for $f$?
To be concrete, consider the AGD and preconditioned AGD iterations (Algorithm˜5, Algorithm˜6),
1input initial point $x^{1}=z^{1}$
2 for $k=1,2,\dots$ do
3 $y^{k}=x^{k}+\tfrac{1}{\sqrt{\kappa}+1}(z^{k}-x^{k})$
4 $x^{k+1}=y^{k}-\tfrac{1}{L}\nabla f(y^{k})$
5 $z^{k+1}=(1-\tfrac{1}{\sqrt{\kappa}})z^{k}+\tfrac{1}{\sqrt{\kappa}}(y^{k}-\tfrac{1}{\mu}\nabla f(y^{k}))$
6 end for
Algorithm 5 Accelerated GD
1input initial point $x^{1}=z^{1}$
2 for $k=1,2,\dots$ do
3 $y^{k}=x^{k}+\tfrac{1}{\sqrt{\kappa}+1}(z^{k}-x^{k})$
4 $x^{k+1}=y^{k}-{P}\nabla f(y^{k})$
5 $z^{k+1}=(1-\tfrac{1}{\sqrt{\kappa}})z^{k}+\tfrac{1}{\sqrt{\kappa}}(y^{k}-\kappa{P}\nabla f(y^{k}))$
6 end for
Algorithm 6 Preconditioned AGD
where the standard $x$ update in Algorithm˜5 is replaced by $x^{k+1}=y^{k}-P_{k}\nabla f(y^{k})$. We believe that the biggest challenge in combining OSGM with AGD lies in feedback design: as in the heavy-ball method, AGD does not guarantee monotonicity in $f(x)$. To faithfully reflect the quality of a stepsize $P$, we likely require another potential function to measure the quality of a stepsize $P$. Unfortunately, potential functions from the literature are all impossible to evaluate efficiently. In fact, they all require knowing the optimum $x^{\star}$ in advance [8]! For example, $\phi(x,z)\coloneqq f(x)-f^{\star}+\tfrac{\mu}{2}\|z-x^{\star}\|^{2}$ is a valid potential function [8]:
$f(x^{k+1})-f^{\star}+\tfrac{\mu}{2}\|z^{k+1}-x^{\star}\|^{2}]\leq(1-\tfrac{1}{\sqrt{\kappa}})[f(x^{k})-f^{\star}+\tfrac{\mu}{2}\|z^{k}-x^{\star}\|^{2}].$
In view of this problem, the challenge is to design a potential function for AGD that does not depend on $x^{\star}$. For quadratic functions, such a potential function exists (Theorem˜5.1).
Theorem 5.1 (Potential function for AGD on quadratic).
Suppose $f$ is a strongly-convex quadratic, and that the iterates $\{(x^{k},z^{k})\}$ are generated by AGD with a fixed value of the momentum $\beta$. Define the potential
$\varphi_{\mu}(x,z)\coloneqq f(z)-f^{\star}+\tfrac{1}{2\mu}\|\nabla f(x)\|^{2}.$
Then
$\varphi_{\mu}(x^{k+1},z^{k+1})\leq\Big[\tfrac{(\kappa-1)^{2}}{\kappa(\sqrt{\kappa}+1)^{2}}(1+\tfrac{1}{2\kappa}+\tfrac{\sqrt{4\kappa+1}}{2\kappa})\Big]~\varphi_{\mu}(x^{k},z^{k})\leq(1-\tfrac{1}{\sqrt{\kappa}})\varphi_{\mu}(x^{k},z^{k}).$
This potential still requires knowledge of $f^{\star}$, but as we have seen in Section˜5.2, it is often possible to relax this requirement as well. Then following the same techniques as OSGM-R [14], we can define ratio feedback $r_{x,z}(P)\coloneqq\tfrac{\varphi_{\mu}(x^{+}(P),z^{+}(P))}{\varphi_{\mu}(x,z)}$, where $x^{+}(P)$ and $z^{+}(P)$ are the iterates after one iteration of Algorithm˜6. Define $\hat{\kappa}$ such that
$1-\tfrac{1}{\sqrt{\hat{\kappa}}}=\min_{P}\max_{x,z}~r_{x,z}(P),$
where $\hat{\kappa}\leq\kappa$ since $r_{x,z}(\frac{1}{L})\leq 1-\frac{1}{\sqrt{\kappa}}$ for any $x,z$. Then an $\mathcal{O}(\sqrt{\hat{\kappa}}\log(1/\varepsilon))$ complexity can be obtained. However, whether it is possible to design an algorithm that achieves $\mathcal{O}(\sqrt{\kappa^{\star}}\log(1/\varepsilon))$ with only blackbox access to $f$ remains unclear.
Feedback design through performance estimation.
OSGM provides a general mechanism to accelerate iterative methods when a computable potential (Lyapunov) function is known. Although such potential functions can often be obtained from literature (for example, Section˜2 demonstrates this approach in the case of heavy-ball momentum), it can be challenging to find a computable potential function that does not rely on $x^{\star}$.
Question 4.
Is there a systematic way to design computable potential functions for iterative methods (that do not rely on the exact solution $x^{\star}$)?
Recent advances in first-order methods have shown that finding a potential function can be reduced to a small-scale semidefinite optimization problem through performance estimation (PEP) [34, 35]. The advantage of the PEP framework is that it can numerically find a tight worst-case potential function for some problem class (e.g., for some given smoothness constant $L$ and strongly convexity constant $\mu$). If PEP can find a potential function that does not depend on $x^{\star}$, then OSGM can be directly applied since it only requires a numerically computable potential function. It is appealing to accelerate first-order methods in practice by applying OSGM to the potential functions generated by performance estimation.
5.4 Convex composite optimization and proximal gradient
OSGM works primarily in the smooth unconstrained setting. Section˜3 shows one way to extend OSGM to smooth nonconvex optimization. Here, we discuss challenges and opportunities for OSGM in nonsmooth optimization.
Question 5.
Can OSGM be used to accelerate the solution of convex composite optimization or general nonsmooth convex optimization problems?
For example, consider the convex composite optimization problem $\min_{x}\varphi(x)\coloneqq f(x)+w(x)$, where $f$ is smooth convex and $w$ is possibly nonsmooth but prox-friendly: that is, for any $\gamma>0$, the proximal mapping
$\operatorname{prox}_{w/\gamma}(x)\coloneqq\operatornamewithlimits{arg\,min}_{z}~\{w(z)+\tfrac{\gamma}{2}\|z-x\|^{2}\}$
can be efficiently evaluated. The operator $\mathcal{G}_{\gamma}(x)\coloneqq\gamma(x-\operatorname{prox}_{w/\gamma}(x-\tfrac{1}{\gamma}\nabla f(x)))$, known as the gradient map, generalizes gradient to this composite setting, and the parameter $\gamma>0$ generalizes the stepsize. The proximal gradient method updates the iterate as
$x^{k+1}=x^{k}-\tfrac{1}{\gamma}\mathcal{G}_{\gamma}(x^{k}).$
One natural idea to accelerate convergence is to adapt the parameter $\gamma$ using OSGM. However, a basic implementation of this idea does not work, as the ratio or hypergradient feedback is not a convex function of $\gamma$ (or $1/\gamma$) due to the complex dependence from $\mathcal{G}_{\gamma}(x)$. A partial solution in this context separates the stepsize multiplying the gradient map from the parameter inside the map: fix $\gamma=L$ in the gradient map and consider the update
$x^{k+1}=x^{k}-P_{k}\mathcal{G}_{L}(x^{k})$
with stepsize $P_{k}$. This iteration directly corresponds to gradient descent. We can show an improved convergence rate in the following two settings:
Case 1.
If $f(x)=0$, then $\mathcal{G}_{L}(x)$ is the gradient of the Moreau envelope $\varphi^{1/L}(x)=w^{1/L}(x)\coloneqq\min_{z}\{w(z)+\tfrac{L}{2}\|z-x\|^{2}\}$, which is smooth and convex. The update is essentially gradient descent on the envelope $w^{1/L}$, and all the OSGM results apply. If $w$ is not prox-friendly, we can solve the proximal subproblem inexactly with an additional subroutine and accelerate gradient descent on the envelope.
Case 2.
When $f$ is a general smooth convex function, we can define the proximal hypergradient feedback
$h_{x}(P)\coloneqq\tfrac{\varphi(x-P\mathcal{G}_{L}(x))-\varphi(x)}{\|\mathcal{G}_{L}(x)\|^{2}}=\tfrac{f(x-P\mathcal{G}_{L}(x))+w(x-P\mathcal{G}_{L}(x))-[f(x)+w(x)]}{\|\mathcal{G}_{L}(x)\|^{2}}.$
Since $w$ is nonsmooth, $h_{x}(P)$ will not be Lipschitz continuous when $\|\mathcal{G}_{L}(x)\|\rightarrow 0$ and online subgradient method cannot guarantee sublinear regret. To address this issue, we consider online proximal gradient update
$P_{k+1}=\operatornamewithlimits{arg\,min}_{P}~\Big\{\tfrac{f(x^{k}-P_{k}\mathcal{G}_{L}(x^{k}))+\langle\nabla f(x^{k}-P_{k}\mathcal{G}_{L}(x^{k}))\mathcal{G}_{L}(x^{k})^{\top},\,P-P_{k}\rangle}{\|\mathcal{G}_{L}(x^{k})\|^{2}}+\tfrac{w(x^{k}-P\mathcal{G}_{L}(x^{k}))}{\|\mathcal{G}_{L}(x^{k})\|^{2}}+\tfrac{1}{2\eta}\|P-P_{k}\|_{F}^{2}\Big\},$
where $h_{x}(P)$ has been partially linearized. This subproblem is a proximal update in $P$ with respect to the (scaled, translated) function $\tilde{w}(P)=w(x-P\mathcal{G}_{L}(x))$. This function often, but not always, inherits from $w$ a proximal mapping that can be efficiently evaluated. If $P$ is a diagonal matrix and $w$ is separable in the coordinates, the proximal operator of $\tilde{w}$ can be evaluated with the same complexity as the proximal operator of $w$. Combined with the monotone landscape action, we can obtain a convergence guarantee in this case in terms of $\|\mathcal{G}_{L}(x)\|$: $\min_{1\leq k\leq K}\|\mathcal{G}_{L}(x^{k})\|^{2}=\mathcal{O}(\tfrac{1}{K})$.
Theorem 5.2 (Informal).
Suppose $f$ is $L$-smooth convex and $w$ is convex. Then given any benchmark stepsize $\hat{P}\in\mathcal{P}$, a variant of OSGM achieves
$\min_{1\leq k\leq K}\|\mathcal{G}_{L}(x^{k})\|^{2}\leq\tfrac{\varphi(x^{1})-\varphi(x^{\star})}{K}\tfrac{1}{\frac{1}{K}\max\{\sum_{k=1}^{K}h_{x^{k}}(\hat{P})-\frac{L}{2}\|P_{1}-\hat{P}\|_{F}^{2},0\}}.$
In particular, taking $P_{1}=\hat{P}=\frac{1}{L}I$ yields $\min_{1\leq k\leq K}\|\mathcal{G}_{L}(x^{k})\|^{2}\leq\tfrac{2L[\varphi(x^{1})-\varphi^{\star}]}{K}$.
A detailed analysis is available in the appendix.
6 Conclusions
The second part of our paper explores the practical aspects of OSGM and outlines several promising directions for future research. We design and analyze OSGM-Best, a practical variant of OSGM that integrates insights into algorithmic behavior with heavy-ball momentum. We also extend the OSGM framework to smooth nonconvex optimization problems. Numerical experiments demonstrate that OSGM achieves competitive performance across a range of real-world optimization tasks.
Appendix
Appendix A Proof of results in Section˜2
A.1 Proof of Lemma 2.1
The assumption $|x_{1}^{k}|\geq\sqrt{2}\kappa^{3/2}|x_{2}^{k}|$ and that $\kappa\geq 2$ implies
$(x_{1}^{k})^{2}\geq 2\kappa^{3}(x_{2}^{k})^{2}\geq 2\kappa^{2}(\kappa-2)(x_{2}^{k})^{2}.$ (18)
Then we deduce that
$\tfrac{\alpha_{k+1}-\alpha_{k}}{\eta}={}$ $\tfrac{(x_{1}^{k})^{2}(1-\alpha_{k})+\kappa^{2}(1-\kappa\alpha_{k})(x_{2}^{k})^{2}}{(x_{1}^{k})^{2}+\kappa^{2}(x_{2}^{k})^{2}}$ (by (3))
$\geq{}$ $\tfrac{1}{4}\tfrac{2(x_{1}^{k})^{2}-2\kappa^{2}(\kappa-2)(x_{2}^{k})^{2}}{(x_{1}^{k})^{2}+\kappa^{2}(x_{2}^{k})^{2}}$ (by $\alpha_{k}\leq\tfrac{1}{2}$)
$\geq{}$ $\tfrac{1}{4}\tfrac{(x_{1}^{k})^{2}}{(x_{1}^{k})^{2}+\kappa^{2}(x_{2}^{k})^{2}}\geq\tfrac{1}{4}.$ (by (18))
This completes the proof.
A.2 Proof of Theorem 2.1
The proof of Theorem˜2.1 requires several lemmas. The proof of each lemma is deferred to the end of the section.
Lemma A.1.
[…]
$J_{{{\texttt{HDM}}}}\left(\alpha^{\star},{\hat{z}^{\star}_{i}}\right)\sim\operatorname{Diag}(1-\alpha^{\star}\lambda_{2},\ldots,1-\alpha^{\star}\lambda_{n-1},J_{{{\texttt{HDM}}}}^{i})\quad\text{and}\quad J_{\texttt{OSGM}}\left(\alpha^{\star},{\hat{z}^{\star}_{i}}\right)\sim\operatorname{Diag}(1-\alpha^{\star}\lambda_{2},\ldots,1-\alpha^{\star}\lambda_{n-1},J_{\texttt{OSGM}}^{i}),$
where $J_{{{\texttt{HDM}}}}^{i}$ and $J_{\texttt{OSGM}}^{i}$, $i=1,2$, are $3\times 3$ matrices defined by
$J_{{{\texttt{HDM}}}}^{1}\coloneqq D\left(\begin{smallmatrix}\frac{\kappa-1}{2\kappa}&-\sqrt{\kappa^{2}+1}&\pm\frac{\sqrt{\kappa^{2}+1}}{\kappa}\\ \frac{-\kappa(\kappa+1)}{(\kappa^{2}+1)^{3/2}}&\frac{\kappa}{\kappa^{2}+1}&\frac{\pm 1}{\kappa^{2}+1}\\ \frac{\mp(\kappa+1)}{(\kappa^{2}+1)^{3/2}}&\frac{\mp 1}{\kappa^{2}+1}&\frac{-1}{\kappa(\kappa^{2}+1)}\end{smallmatrix}\right)D^{-1},$ $\quad J_{{{\texttt{HDM}}}}^{2}\coloneqq D\left(\begin{smallmatrix}\frac{\kappa-1}{2\kappa}&\sqrt{\kappa^{2}+1}&\pm\frac{\sqrt{\kappa^{2}+1}}{\kappa}\\ \frac{\kappa(\kappa+1)}{(\kappa^{2}+1)^{3/2}}&\frac{\kappa}{\kappa^{2}+1}&\frac{\mp 1}{\kappa^{2}+1}\\ \frac{\mp(\kappa+1)}{(\kappa^{2}+1)^{3/2}}&\frac{\pm 1}{\kappa^{2}+1}&\frac{-1}{\kappa(\kappa^{2}+1)}\end{smallmatrix}\right)D^{-1};$
$J_{\texttt{OSGM}}^{1}\coloneqq D\left(\begin{smallmatrix}\frac{\kappa-1}{2\kappa}&-\sqrt{\kappa^{2}+1}&\pm\frac{\sqrt{\kappa^{2}+1}}{\kappa}\\ \frac{-\kappa(\kappa+1)}{(\kappa^{2}+1)^{3/2}}&\frac{-\kappa^{2}}{\kappa^{2}+1}&\frac{\pm\kappa}{\kappa^{2}+1}\\ \frac{\mp(\kappa+1)}{(\kappa^{2}+1)^{3/2}}&\frac{\mp\kappa}{\kappa^{2}+1}&\frac{1}{\kappa^{2}+1}\end{smallmatrix}\right)D^{-1},$ $\quad J_{\texttt{OSGM}}^{2}\coloneqq D\left(\begin{smallmatrix}\frac{\kappa-1}{2\kappa}&\sqrt{\kappa^{2}+1}&\pm\frac{\sqrt{\kappa^{2}+1}}{\kappa}\\ \frac{\kappa(\kappa+1)}{(\kappa^{2}+1)^{3/2}}&\frac{-\kappa^{2}}{\kappa^{2}+1}&\frac{\mp\kappa}{\kappa^{2}+1}\\ \frac{\mp(\kappa+1)}{(\kappa^{2}+1)^{3/2}}&\frac{\pm\kappa}{\kappa^{2}+1}&\frac{1}{\kappa^{2}+1}\end{smallmatrix}\right)D^{-1},$
and $D=\mathrm{Diag}(\tfrac{\kappa-1}{L(\kappa+1)},1,1)$ is a $3\times 3$ invertible diagonal matrix.
The stability of the period-$2$ orbit is determined by the spectral radius of the products $J_{{{\texttt{HDM}}}}\left(\alpha^{\star},{\hat{z}^{\star}_{1}}\right)J_{{{\texttt{HDM}}}}\left(\alpha^{\star},{\hat{z}^{\star}_{2}}\right)$ and $J_{\texttt{OSGM}}\left(\alpha^{\star},{\hat{z}^{\star}_{1}}\right)J_{\texttt{OSGM}}\left(\alpha^{\star},{\hat{z}^{\star}_{2}}\right)$, which equals the maximum absolute value of the eigenvalues. By Lemma˜A.1, the spectrum of these two matrices are
$\sigma\left(J_{{{\texttt{HDM}}}}\left(\alpha^{\star},{\hat{z}^{\star}_{1}}\right)J_{{{\texttt{HDM}}}}\left(\alpha^{\star},{\hat{z}^{\star}_{2}}\right)\right)$ $=\{(1-\alpha^{\star}\lambda_{2})^{2},\ldots,(1-\alpha^{\star}\lambda_{n-1})^{2}\}\cup\sigma(J_{{{\texttt{HDM}}}}^{1}J_{{{\texttt{HDM}}}}^{2});$
$\sigma\left(J_{\texttt{OSGM}}\left(\alpha^{\star},{\hat{z}^{\star}_{1}}\right)J_{\texttt{OSGM}}\left(\alpha^{\star},{\hat{z}^{\star}_{2}}\right)\right)$ $=\{(1-\alpha^{\star}\lambda_{2})^{2},\ldots,(1-\alpha^{\star}\lambda_{n-1})^{2}\}\cup\sigma(J_{\texttt{OSGM}}^{1}J_{\texttt{OSGM}}^{2}).$
Since the eigenvalues of $A$ are distinct, i.e., $L=\lambda_{1}>\lambda_{2}>\cdots>\lambda_{n-1}>\lambda_{n}=\mu$, and $\alpha^{\star}=\tfrac{2}{L+\mu}$, the first n2n-2 eigenvalues satisfy
$\tfrac{\kappa-1}{\kappa+1}=1-\alpha^{\star}L<1-\alpha^{\star}\lambda_{i}<1-\alpha^{\star}\mu=\tfrac{\kappa-1}{\kappa+1}\quad\text{for }i=2,\ldots,n-1.$
Then $(1-\alpha^{\star}\lambda_{i})^{2}<1$ for all $i=2,\ldots,n-1$ and the stability of the orbit $\{(\alpha^{\star},\hat{z}^{\star}_{1}),(\alpha^{\star},\hat{z}^{\star}_{2})\}$ is determined by the spectral radius of $J_{{{\texttt{HDM}}}}^{1}J_{{{\texttt{HDM}}}}^{2}$ and $J_{\texttt{OSGM}}^{1}J_{\texttt{OSGM}}^{2}$ for the corresponding system.
Lemma A.2.
The $3\times 3$ matrices $J_{{{\texttt{HDM}}}}^{1}J_{{{\texttt{HDM}}}}^{2}$ and $J_{\texttt{OSGM}}^{1}J_{\texttt{OSGM}}^{2}$ have spectral radius
$\rho(J_{{{\texttt{HDM}}}}^{1}J_{{{\texttt{HDM}}}}^{2})={}$ $\max\Big\{-\tfrac{3\kappa^{3}-7\kappa-9}{8(\kappa+1)\kappa^{2}}\pm\Big[\tfrac{\sqrt{-(\kappa-3)^{2}\kappa^{2}(\kappa+1)^{3}(7\kappa-9)(\kappa^{2}+1)^{5}}}{8(\kappa^{2}+1)^{5/2}(\kappa+1)\kappa^{3}}-\tfrac{5}{8(\kappa+1)}\Big]\Big\}\leq 1$
$\rho(J_{\texttt{OSGM}}^{1}J_{\texttt{OSGM}}^{2})={}$ $\tfrac{3\kappa+1}{2\kappa}>1$
Observe that the spectral radius $\rho(J_{{{\texttt{HDM}}}}^{1}J_{{{\texttt{HDM}}}}^{2})$ asymptotically increases to $\tfrac{1}{2}$ when $\kappa\rightarrow\infty$, implying that the eigenvalues of the $3\times 3$ submatrix $J_{{{\texttt{HDM}}}}^{1}J_{{{\texttt{HDM}}}}^{2}$ are not active for the spectral radius of the whole Jacobian when $\kappa$ is large. In this case, the spectral radius of $J_{{{\texttt{HDM}}}}\left(\alpha^{\star},{\hat{z}^{\star}_{1}}\right)J_{{{\texttt{HDM}}}}\left(\alpha^{\star},{\hat{z}^{\star}_{2}}\right)$ is the maximum of $(1-\alpha^{\star}\lambda_{2})^{2}$ and $(1-\alpha^{\star}\lambda_{n-1})^{2}$.
Remark 3.
While our result is established for $\eta=\frac{1}{L}$, the same quantitative result holds for $\eta=\frac{\theta}{L},\theta\in(0,1]$.
A.2.1 Proof of Lemma 2.2
To show the orbit, we simply plug in $(\alpha^{\star},\hat{z}^{\star}_{1})$ and $(\alpha^{\star},\hat{z}^{\star}_{2})$ into the updates of both HDM and OSGM. Observe that $L^{2}\gamma^{2}=\mu^{2}\delta^{2}$ and the eigenvectors $q_{1}$ and $q_{n}$ are orthogonal to each other. Then
$\langle\hat{z}^{\star}_{1},\Lambda^{3}\hat{z}^{\star}_{1}\rangle$ $=\langle\gamma e_{1}+\delta e_{n},L^{3}\gamma e_{1}+\mu^{3}\delta e_{n}\rangle=L^{3}\gamma^{2}+\mu^{3}\delta^{2}=(L+\mu)L^{2}\gamma^{2},$ (19)
$\langle\hat{z}^{\star}_{2},\Lambda^{3}\hat{z}^{\star}_{2}\rangle$ $=\langle-\gamma e_{1}+\delta e_{n},-L^{3}\gamma e_{1}+\mu^{3}\delta e_{n}\rangle=L^{3}\gamma^{2}+\mu^{3}\delta^{2}=(L+\mu)L^{2}\gamma^{2},$ (20)
$\langle\hat{z}^{\star}_{1},\Lambda^{2}\hat{z}^{\star}_{1}\rangle$ $=\langle\gamma e_{1}+\delta e_{n},L^{2}\gamma e_{1}+\mu^{2}\delta e_{n}\rangle=L^{2}\gamma^{2}+\mu^{2}\delta^{2}=2L^{2}\gamma^{2},$ (21)
$\langle\hat{z}^{\star}_{2},\Lambda^{2}\hat{z}^{\star}_{2}\rangle$ $=\langle-\gamma e_{1}+\delta e_{n},-L^{2}\gamma e_{1}+\mu^{2}\delta e_{n}\rangle=L^{2}\gamma^{2}+\mu^{2}\delta^{2}=2L^{2}\gamma^{2},$ (22)
and hence
$\tfrac{\langle\hat{z}^{\star}_{1},\Lambda^{3}\hat{z}^{\star}_{1}\rangle}{\langle\hat{z}^{\star}_{1},\Lambda^{2}\hat{z}^{\star}_{1}\rangle}=\tfrac{\langle\hat{z}^{\star}_{2},\Lambda^{3}\hat{z}^{\star}_{2}\rangle}{\langle\hat{z}^{\star}_{2},\Lambda^{2}\hat{z}^{\star}_{2}\rangle}=\tfrac{L+\mu}{2}.$ (23)
The $\alpha$-update of both systems $\mathcal{F}_{{{\texttt{HDM}}}}$} and $\mathcal{F}_{{{\texttt{OSGM}}}}$} are the same. Starting from either state $(\alpha^{\star},\hat{z}^{\star}_{1})$ and $(\alpha^{\star},\hat{z}^{\star}_{2})$, the new state of stepsize is
$\alpha^{\star}]^{+}(\hat{z}^{\star}_{i})=\alpha^{\star}+\eta-\alpha^{\star}\eta\tfrac{\langle\hat{z}^{\star}_{i},\Lambda^{3}\hat{z}^{\star}_{i}\rangle}{\langle\hat{z}^{\star}_{i},\Lambda^{2}\hat{z}^{\star}_{i}\rangle}=\alpha^{\star}+\eta-\alpha^{\star}\eta\tfrac{L+\mu}{2}=\alpha^{\star}.$ (24)
Since the new state of stepsize remains unchanged to be $\alpha^{\star}=\tfrac{L+\mu}{2}$, both systems end up with the same new state:
${\mathcal{F}_{{{\texttt{HDM}}}}}(\alpha^{\star},\hat{z}^{\star}_{i})$ $=([\alpha^{\star}]^{+}(\hat{z}^{\star}_{i}),\tfrac{(I-[\alpha^{\star}]^{+}(\hat{z}^{\star}_{i})\Lambda)\hat{z}^{\star}_{i}}{\|(I-[\alpha^{\star}]^{+}(\hat{z}^{\star}_{i})\Lambda)\hat{z}^{\star}_{i}\|})=(\alpha^{\star},\tfrac{(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{i}}{\|(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{i}\|});$
${\mathcal{F}_{{{\texttt{OSGM}}}}}(\alpha^{\star},\hat{z}^{\star}_{i})$ $=([\alpha^{\star}]^{+}(\hat{z}^{\star}_{i}),\tfrac{(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{i}}{\|(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{i}\|})=(\alpha^{\star},\tfrac{(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{i}}{\|(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{i}\|}).$
The updates of direction before the normalization are
$(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{1}$ $=(1-\tfrac{2L}{L+\mu})\gamma e_{1}+(1-\tfrac{2\mu}{L+\mu})\delta e_{n}=\tfrac{L-\mu}{L+\mu}(-\gamma e_{1}+\delta e_{n})=\tfrac{\kappa-1}{\kappa+1}\hat{z}^{\star}_{2};$ (25)
$(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{2}$ $=-(1-\tfrac{2L}{L+\mu})\gamma e_{1}+(1-\tfrac{2\mu}{L+\mu})\delta e_{n}=\tfrac{L-\mu}{L+\mu}(\gamma e_{1}+\delta e_{n})=\tfrac{\kappa-1}{\kappa+1}\hat{z}^{\star}_{1}.$ (26)
Since $\hat{x}^{\star}_{1}$ and $\hat{x}^{\star}_{2}$ both have unit norm, after the normalization, we have
${\mathcal{F}_{{{\texttt{HDM}}}}}(\alpha^{\star},\hat{z}^{\star}_{1})={\mathcal{F}_{{{\texttt{OSGM}}}}}(\alpha^{\star},\hat{z}^{\star}_{1})=(\alpha^{\star},\hat{z}^{\star}_{2})\quad\text{and}\quad{\mathcal{F}_{{{\texttt{HDM}}}}}(\alpha^{\star},\hat{z}^{\star}_{2})={\mathcal{F}_{{{\texttt{OSGM}}}}}(\alpha^{\star},\hat{z}^{\star}_{2})=(\alpha^{\star},\hat{z}^{\star}_{1}).$
This proves that $\{(\alpha^{\star},\hat{z}^{\star}_{1}),(\alpha^{\star},\hat{z}^{\star}_{2})\}$ is an orbit with period $2$.
A.2.2 Proof of Lemma A.1
The Jacobians of $\mathcal{F}_{{{\texttt{HDM}}}}$} and $\mathcal{F}_{{{\texttt{OSGM}}}}$} are
[…]
where $\Pi(\alpha,\hat{z})\coloneqq I-\tfrac{u(\alpha,\hat{z})u(\alpha,\hat{z})^{\top}}{\|u(\alpha,\hat{z})\|^{2}}$, $u(\alpha,\hat{x})\coloneqq(I-\alpha\Lambda)\hat{z}$, and $v(\alpha,\hat{z})\coloneqq\tfrac{2\alpha\eta}{\langle\hat{z},\Lambda^{2}\hat{z}\rangle}\Lambda^{2}(\tfrac{\langle\hat{z},\Lambda^{3}\hat{z}\rangle}{\langle\hat{z},\Lambda^{2}\hat{z}\rangle}I-\Lambda)\hat{z}$. To evaluate the Jacobians at $(\alpha^{\star},\hat{z}^{\star}_{1})$ and $(\alpha^{\star},\hat{z}^{\star}_{2})$, we calculate each block one by one. Note that $J_{{{\texttt{HDM}}}}$ and $J_{\texttt{OSGM}}$ only differ in the last row.
$(1,1)$-block of $J_{{{\texttt{HDM}}}}(\alpha^{\star},\hat{z}^{\star}_{i})$ and $J_{\texttt{OSGM}}(\alpha^{\star},\hat{z}^{\star}_{i})$.
Equation (23) and $\eta=\tfrac{1}{L}$ imply
$1-\eta\tfrac{\langle\hat{z}^{\star}_{i},\Lambda^{3}\hat{z}^{\star}_{i}\rangle}{\langle\hat{z}^{\star}_{i},\Lambda^{2}\hat{z}^{\star}_{i}\rangle}=1-\tfrac{L+\mu}{2L}=\tfrac{\kappa-1}{2\kappa}\quad\text{for }i=1,2.$ (33)
$(1,2)$-block of $J_{{{\texttt{HDM}}}}(\alpha^{\star},\hat{z}^{\star}_{i})$ and $J_{\texttt{OSGM}}(\alpha^{\star},\hat{z}^{\star}_{i})$.
Using (21)–(23) and $\eta=\tfrac{1}{L}$, we have
$v(\alpha^{\star},\hat{z}^{\star}_{i})=\tfrac{2\alpha^{\star}\eta}{\langle\hat{z}^{\star}_{i},\Lambda^{2}\hat{z}^{\star}_{i}\rangle}\Lambda^{2}(\tfrac{\langle\hat{z}^{\star}_{i},\Lambda^{3}\hat{z}^{\star}_{i}\rangle}{\langle\hat{z}^{\star}_{i},\Lambda^{2}\hat{z}^{\star}_{i}\rangle}I-\Lambda)\hat{z}^{\star}_{i}=\tfrac{1}{L^{3}\gamma^{2}}\Lambda^{2}(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{i},$
which, together with (25) and (26), implies
$v(\alpha^{\star},\hat{z}^{\star}_{1})$ $=\tfrac{\kappa-1}{L^{3}\gamma^{2}(\kappa+1)}\Lambda^{2}\hat{z}^{\star}_{2}=\tfrac{\kappa-1}{L^{3}\gamma^{2}(\kappa+1)}[-L^{2}\gamma e_{1}+\mu^{2}\delta e_{n}]=-\tfrac{(\kappa-1)\sqrt{\kappa^{2}+1}}{L(\kappa+1)}e_{1}+\tfrac{\pm(\kappa-1)\sqrt{\kappa^{2}+1}}{L\kappa(\kappa+1)}e_{n};$ (34)
$v(\alpha^{\star},\hat{z}^{\star}_{2})$ $=\tfrac{\kappa-1}{L^{3}\gamma^{2}(\kappa+1)}\Lambda^{2}\hat{z}^{\star}_{1}=\tfrac{\kappa-1}{L^{3}\gamma^{2}(\kappa+1)}[L^{2}\gamma e_{1}+\mu^{2}\delta e_{n}]=\tfrac{(\kappa-1)\sqrt{\kappa^{2}+1}}{L(\kappa+1)}e_{1}+\tfrac{\pm(\kappa-1)\sqrt{\kappa^{2}+1}}{L\kappa(\kappa+1)}e_{n}.$ (35)
$(2,1)$-block of $J_{\texttt{OSGM}}(\alpha^{\star},\hat{z}^{\star}_{i})$.
Equations (25)–(26) imply
$u(\alpha^{\star},\hat{z}^{\star}_{1})=(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{1}=\tfrac{\kappa-1}{\kappa+1}\hat{z}^{\star}_{2},\quad u(\alpha^{\star},\hat{z}^{\star}_{2})=(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{2}=\tfrac{\kappa-1}{\kappa+1}\hat{z}^{\star}_{1},\quad\|u(\alpha^{\star},\hat{z}^{\star}_{1})\|=\|u(\alpha^{\star},\hat{z}^{\star}_{2})\|=\tfrac{\kappa-1}{\kappa+1},$ (36)
and hence
$I-\tfrac{u(\alpha^{\star},\hat{z}^{\star}_{1})u(\alpha^{\star},\hat{z}^{\star}_{1})^{\top}}{\|u(\alpha^{\star},\hat{z}^{\star}_{1})\|^{2}}=I-\hat{z}^{\star}_{2}{(\hat{z}^{\star}_{2})}^{\top}\quad\text{and}\quad I-\tfrac{u(\alpha^{\star},\hat{z}^{\star}_{2})u(\alpha^{\star},\hat{z}^{\star}_{2})^{\top}}{\|u(\alpha^{\star},\hat{z}^{\star}_{2})\|^{2}}=I-\hat{z}^{\star}_{1}{(\hat{z}^{\star}_{1})}^{\top}.$ (37)
Observe that ${(\hat{z}^{\star}_{2})}^{\top}A\hat{z}^{\star}_{1}={(\hat{z}^{\star}_{1})}^{\top}A\hat{z}^{\star}_{2}=-L\gamma^{2}+\mu\delta^{2}$, together with (36), we have
$\tfrac{-1}{\|u(\alpha^{\star},\hat{z}^{\star}_{1})\|}\Pi(\alpha^{\star},\hat{z}^{\star}_{1})\Lambda\hat{z}^{\star}_{1}$ $=\tfrac{-1}{\|u(\alpha^{\star},\hat{z}^{\star}_{1})\|}(I-\tfrac{u(\alpha^{\star},\hat{z}^{\star}_{1})u(\alpha^{\star},\hat{z}^{\star}_{1})^{\top}}{\|u(\alpha^{\star},\hat{z}^{\star}_{1})\|^{2}})\Lambda\hat{z}^{\star}_{1}$
$=-\tfrac{\kappa+1}{\kappa-1}(\Lambda\hat{z}^{\star}_{1}-\hat{z}^{\star}_{2}{(\hat{z}^{\star}_{2})}^{\top}\Lambda\hat{z}^{\star}_{1})$
$=-\tfrac{\kappa+1}{\kappa-1}[\Lambda\hat{z}^{\star}_{1}-(-L\gamma^{2}+\mu\delta^{2})\hat{z}^{\star}_{2}]$
$=-\tfrac{\kappa+1}{\kappa-1}[(L\gamma e_{1}+\mu\delta e_{n})-(-L\gamma^{2}+\mu\delta^{2})(-\gamma e_{1}+\delta e_{n})]$
$=-\tfrac{\kappa+1}{\kappa-1}[(L\gamma-L\gamma^{3}+\mu\gamma\delta^{2})e_{1}+(\mu\delta+L\gamma^{2}\delta-\mu\delta^{3})e_{n}]$
$=\tfrac{-L\kappa(\kappa+1)^{2}}{(\kappa-1)(\kappa^{2}+1)^{3/2}}e_{1}\mp\tfrac{L(\kappa+1)^{2}}{(\kappa-1)(\kappa^{2}+1)^{3/2}}e_{n}$ (38)
$\tfrac{-1}{\|u(\alpha^{\star},\hat{z}^{\star}_{1})\|}\Pi(\alpha^{\star},\hat{z}^{\star}_{2})\Lambda\hat{z}^{\star}_{2}$ $=\tfrac{-1}{\|u(\alpha^{\star},\hat{z}^{\star}_{2})\|}(I-\tfrac{u(\alpha^{\star},\hat{z}^{\star}_{2})u(\alpha^{\star},\hat{z}^{\star}_{2})^{\top}}{\|u(\alpha^{\star},\hat{z}^{\star}_{2})\|^{2}})\Lambda\hat{z}^{\star}_{2}$
$=-\tfrac{\kappa+1}{\kappa-1}(\Lambda\hat{z}^{\star}_{2}-\hat{z}^{\star}_{1}{(\hat{z}^{\star}_{1})}^{\top}\Lambda\hat{z}^{\star}_{2})$
$=-\tfrac{\kappa+1}{\kappa-1}[\Lambda\hat{z}^{\star}_{2}-(-L\gamma^{2}+\mu\delta^{2})\hat{z}^{\star}_{1}]$
$=-\tfrac{\kappa+1}{\kappa-1}[(-L\gamma e_{1}+\mu\delta e_{n})-(-L\gamma^{2}+\mu\delta^{2})(\gamma e_{1}+\delta e_{n})]$
$=-\tfrac{\kappa+1}{\kappa-1}[(-L\gamma+L\gamma^{3}-\mu\gamma\delta^{2})e_{1}+(\mu\delta+L\gamma^{2}\delta-\mu\delta^{3})e_{n}]$
$=\tfrac{L\kappa(\kappa+1)^{2}}{(\kappa-1)(\kappa^{2}+1)^{3/2}}e_{1}\mp\tfrac{L(\kappa+1)^{2}}{(\kappa-1)(\kappa^{2}+1)^{3/2}}e_{n}.$ (39)
$(2,1)$-block of $J_{{\texttt{HDM}}}(\alpha^{\star},\hat{z}^{\star}_{i})$.
Note that the $(2,1)$-block of $J_{{\texttt{HDM}}}(\alpha^{\star},\hat{z}^{\star}_{i})$ is a scalar multiple of the $(2,1)$-block of $J_{\texttt{OSGM}}(\alpha^{\star},\hat{z}^{\star}_{i})$ with the scalar $1-\eta\tfrac{\langle\hat{z}^{\star}_{i},\Lambda^{3}\hat{z}^{\star}_{i}\rangle}{\langle\hat{z}^{\star}_{i},\Lambda^{2}\hat{z}^{\star}_{i}\rangle}$. Then using (33) and (38)–(39), we have
$\tfrac{-1}{\|u(\alpha^{\star},\hat{z}^{\star}_{1})\|}\Pi(\alpha^{\star},\hat{z}^{\star}_{1})\big(1-\eta\tfrac{\langle\hat{z}^{\star}_{1},\Lambda^{3}\hat{z}^{\star}_{1}\rangle}{\langle\hat{z}^{\star}_{1},\Lambda^{2}\hat{z}^{\star}_{1}\rangle}\big)\Lambda\hat{z}^{\star}_{1}$ $=\tfrac{-L\kappa(\kappa+1)^{2}}{2\kappa(\kappa^{2}+1)^{3/2}}e_{1}\mp\tfrac{L(\kappa+1)^{2}}{2\kappa(\kappa^{2}+1)^{3/2}}e_{n};$
$\tfrac{-1}{\|u(\alpha^{\star},\hat{z}^{\star}_{2})\|}\Pi(\alpha^{\star},\hat{z}^{\star}_{2})\big(1-\eta\tfrac{\langle\hat{z}^{\star}_{2},\Lambda^{3}\hat{z}^{\star}_{2}\rangle}{\langle\hat{z}^{\star}_{2},\Lambda^{2}\hat{z}^{\star}_{2}\rangle}\big)\Lambda\hat{z}^{\star}_{2}$ $=\tfrac{L\kappa(\kappa+1)^{2}}{2\kappa(\kappa^{2}+1)^{3/2}}e_{1}\mp\tfrac{L(\kappa+1)^{2}}{2\kappa(\kappa^{2}+1)^{3/2}}e_{n}.$
$(2,2)$-block of $J_{\texttt{OSGM}}(\alpha^{\star},\hat{z}^{\star}_{i})$.
Let $I_{1,n}\coloneqq[e_{1},~e_{n}]\in\mathbb{R}^{n\times 2}$ denote the first and last column of an $n\times n$ identity matrix. Using (36) and (37), we have
$\tfrac{1}{\|u(\alpha^{\star},\hat{z}^{\star}_{1})\|}\Pi(\alpha^{\star},\hat{z}^{\star}_{1})(I-\alpha^{\star}\Lambda)$ $=\tfrac{\kappa+1}{\kappa-1}\big(I-\hat{z}^{\star}_{2}{(\hat{z}^{\star}_{2})}^{\top}\big)(I-\alpha^{\star}\Lambda)=\tfrac{\kappa+1}{\kappa-1}(I-\alpha^{\star}\Lambda)-\tfrac{\kappa+1}{\kappa-1}\hat{z}^{\star}_{2}[(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{2}]^{\top}$
$=\tfrac{\kappa+1}{\kappa-1}(I-\alpha^{\star}\Lambda)-\hat{z}^{\star}_{2}{(\hat{z}^{\star}_{1})}^{\top}$
$=\tfrac{\kappa+1}{\kappa-1}(I-\alpha^{\star}\Lambda)-I_{1,n}\left(\begin{smallmatrix}-\gamma^{2}&-\gamma\delta\\ \gamma\delta&\delta^{2}\end{smallmatrix}\right)I^{\top}_{1,n},$ (40)
$\tfrac{1}{\|u(\alpha^{\star},\hat{z}^{\star}_{2})\|}\Pi(\alpha^{\star},\hat{z}^{\star}_{2})(I-\alpha^{\star}\Lambda)$ $=\tfrac{\kappa+1}{\kappa-1}\big(I-\hat{z}^{\star}_{1}{(\hat{z}^{\star}_{1})}^{\top}\big)(I-\alpha^{\star}\Lambda)=\tfrac{\kappa+1}{\kappa-1}(I-\alpha^{\star}\Lambda)-\tfrac{\kappa+1}{\kappa-1}\hat{z}^{\star}_{1}[(I-\alpha^{\star}\Lambda)\hat{z}^{\star}_{1}]^{\top}$
$=\tfrac{\kappa+1}{\kappa-1}(I-\alpha^{\star}\Lambda)-\hat{z}^{\star}_{1}{(\hat{z}^{\star}_{2})}^{\top}$
$=\tfrac{\kappa+1}{\kappa-1}(I-\alpha^{\star}\Lambda)-I_{1,n}\left(\begin{smallmatrix}-\gamma^{2}&\gamma\delta\\ -\gamma\delta&\delta^{2}\end{smallmatrix}\right)I^{\top}_{1,n}.$ (41)
Both (40) and (41) have $1-\alpha^{\star}\lambda_{i}$ at the $(i,i)$-entry for $i=2,\ldots,n-1$ and the only other non-zero entries lie on the four corners indexed by $(1,1)$, $(1,n)$, $(n,1)$, and $(n,n)$.
$(2,2)$-block of $J_{{\texttt{HDM}}}(\alpha^{\star},\hat{z}^{\star}_{i})$.
Equation (24) reads as $\alpha^{\star}]^{+}(\hat{z}^{\star}_{i})=\alpha^{\star}$ for $i=1,2$, and thus the $(2,2)$-block of $J_{{{\texttt{HDM}}}}(\alpha^{\star},\hat{z}_{i}^{\star})$ becomes
$\tfrac{1}{\|u(\alpha^{\star},\hat{z}^{\star}_{i})\|}\Pi(\alpha^{\star},\hat{z}^{\star}_{i})[(I-[\alpha^{\star}]^{+}(\hat{z}^{\star}_{i})\Lambda)-\Lambda\hat{z}^{\star}_{i}v(\alpha^{\star},\hat{z}^{\star}_{i})^{\top}]=\tfrac{1}{\|u(\alpha^{\star},\hat{z}^{\star}_{i})\|}\Pi(\alpha^{\star},\hat{z}^{\star}_{i})[(I-\alpha^{\star}\Lambda)-\Lambda\hat{z}^{\star}_{i}v(\alpha^{\star},\hat{z}^{\star}_{i})^{\top}],$
in which is first term has been evaluated in (40) and (41). Using (34)–(35) and (38)–(39), the second term is evaluated as
$-\tfrac{1}{\|u(\alpha^{\star},\hat{z}^{\star}_{1})\|}\Pi(\alpha^{\star},\hat{z}^{\star}_{1})\Lambda\hat{z}^{\star}_{1}v(\alpha^{\star},\hat{z}^{\star}_{1})^{\top}$ $=-\tfrac{1}{L^{3}\gamma^{2}}I_{1,n}\left(\begin{smallmatrix}L\gamma-L\gamma^{3}+\mu\gamma\delta^{2}\\ \mu\delta+L\gamma^{2}\delta-\mu\delta^{3}\end{smallmatrix}\right)\left(\begin{smallmatrix}-L^{2}\gamma\\ \mu^{2}\delta\end{smallmatrix}\right)^{\top}I_{1,n}^{\top};$ (42)
$-\tfrac{1}{\|u(\alpha^{\star},\hat{z}^{\star}_{2})\|}\Pi(\alpha^{\star},\hat{z}^{\star}_{2})\Lambda\hat{z}^{\star}_{2}v(\alpha^{\star},\hat{z}^{\star}_{2})^{\top}$ $=-\tfrac{1}{L^{3}\gamma^{2}}I_{1,n}\left(\begin{smallmatrix}-L\gamma+L\gamma^{3}-\mu\gamma\delta^{2}\\ \mu\delta+L\gamma^{2}\delta-\mu\delta^{3}\end{smallmatrix}\right)\left(\begin{smallmatrix}L^{2}\gamma\\ \mu^{2}\delta\end{smallmatrix}\right)^{\top}I^{\top}_{1,n},$ (43)
The non-zero entries of (42) and (43) again only lie on the four corners indexed by $(1,1)$, $(1,n)$, $(n,1)$, and $(n,n)$. Hence, the Jacobians $J_{{{\texttt{HDM}}}}(\alpha^{\star},\hat{z}^{\star}_{i})$ and $J_{{\texttt{OSGM}}}(\alpha^{\star},\hat{z}^{\star}_{i})$ for $i=1,2$ all have the structure
$\left(\begin{array}[]{ccccccc}\times&\times&0&&\cdots&0&\times\\ \times&\times&0&&\cdots&&\times\\ 0&0&1-\alpha^{\star}\lambda_{2}&&&&0\\ &&&\ddots&&&\\ \vdots&\vdots&&&\ddots&&\vdots\\ 0&0&&&&1-\alpha^{\star}\lambda_{n-1}&0\\ \times&\times&0&&\cdots&0&\times\end{array}\right).$
We can simultaneously permute the rows and columns so that all the four Jacobians are similar to a block diagonal matrix that consists of n2n-2 individual diagonal entries $1-\alpha^{\star}\lambda_{2},\ldots,1-\alpha^{\star}\lambda_{n-1}$ and one $3\times 3$ dense matrix, denoted by $J_{{{\texttt{HDM}}}}^{i}$ and $J_{{\texttt{OSGM}}}^{i}$ for $i=1,2$. After careful but tedious calculations, the four $3\times 3$ matrices match their definition in the lemma.
A.2.3 Proof of Lemma A.2
The result is obtained by direct computation using Mathematica.
A.3 Proof of Lemma 2.3
The proof is adapted from [7, Theorem 2] for completeness. The relation $x^{+}=x-\alpha\nabla f(x)+\beta(x-x^{-})$ implies
$x^{+}-x$ $=-\alpha\nabla f(x)+\beta(x-x^{-}),$ (44)
$\|x^{+}-x\|^{2}$ $=\alpha^{2}\|\nabla f(x)\|^{2}+\beta^{2}\|x-x^{-}\|^{2}-2\alpha\beta\langle\nabla f(x),x-x^{-}\rangle.$ (45)
The $L$-smoothness of $f$ implies
$f(x^{+})\leq f(x)+\langle\nabla f(x),x^{+}-x\rangle+\tfrac{L}{2}\|x^{+}-x\|^{2}.$ (46)
Plugging (44) and (45) into (46) gives
$f(x^{+})$ $\leq f(x)-\alpha\|\nabla f(x)\|^{2}+\beta\langle\nabla f(x),x-x^{-}\rangle$
$\quad+\tfrac{\alpha^{2}L}{2}\|\nabla f(x)\|^{2}+\tfrac{\beta^{2}L}{2}\|x-x^{-}\|^{2}-L\alpha\beta\langle\nabla f(x),x-x^{-}\rangle$
$=f(x)+(\tfrac{\alpha^{2}L}{2}-\alpha)\|\nabla f(x)\|^{2}+\beta(1-L\alpha)\langle\nabla f(x),x-x^{-}\rangle+\tfrac{\beta^{2}L}{2}\|x-x^{-}\|^{2}.$ (47)
Multiplying (45) by $\tfrac{1-\alpha L}{2\alpha}$ and adding to (47), we have
$f(x^{+})+\tfrac{1-\alpha L}{2\alpha}\|x^{+}-x\|^{2}$ $\leq f(x)+(\tfrac{\alpha^{2}L}{2}-\alpha)\|\nabla f(x)\|^{2}+\beta(1-L\alpha)\langle\nabla f(x),x-x^{-}\rangle+\tfrac{\beta^{2}L}{2}\|x-x^{-}\|^{2}$
$\quad+\tfrac{\alpha(1-\alpha L)}{2}\|\nabla f(x)\|^{2}+\tfrac{\beta^{2}(1-\alpha L)}{2\alpha}\|x-x^{-}\|^{2}-\beta(1-L\alpha)\langle\nabla f(x),x-x^{-}\rangle$
$=f(x)+\tfrac{\beta^{2}}{2\alpha}\|x-x^{-}\|^{2}-\tfrac{\alpha}{2}\|\nabla f(x)\|^{2}$
$=f(x)+\tfrac{1-\alpha L}{2\alpha}\|x-x^{-}\|^{2}-\tfrac{\alpha}{2}[\|\nabla f(x)\|^{2}+\tfrac{1-\alpha L-\beta^{2}}{\alpha^{2}}\|x-x^{-}\|^{2}].$ (48)
Let $\omega=\tfrac{1-\alpha L}{2\alpha}$. Then (48) becomes
$\varphi_{\omega}(x^{+},x)\leq\varphi_{\omega}(x,x^{-})-\tfrac{\alpha}{2}\big[\|\nabla f(x)\|^{2}+\tfrac{1-\alpha L-\beta^{2}}{\alpha^{2}}\|x-x^{-}\|^{2}\big]$
and this completes the proof.
A.4 Proof of Lemma 2.4
Notation.
Recall that $z=(x,x^{-})$. We explicitly use $(x,x^{-})$ when establishing the analytic properties of $\varphi_{\omega}$.
Convexity.
To prove the convexity of $h_{x,x^{-}}(P,\beta)$ in $(P,\beta)$, it suffices to show the joint convexity of
$\varphi_{\omega}(x^{+}(P,\beta),x)\coloneqq f(x^{+}(P,\beta))+\tfrac{\omega}{2}\|x^{+}(P,\beta)-x\|^{2}.$
Since $x^{+}(P,\beta)\coloneqq x-P\nabla f(x)+\beta(x-x^{-})$ is affine in $(P,\beta)$ and $f$ and $\|\cdot\|^{2}$ are both convex, the composition rule of convex functions implies the joint convexity of $\varphi_{\omega}(x^{+}(P,\beta),x)$.
Gradient.
The chain rule implies
$\nabla_{P}h_{x,x^{-}}(P,\beta)$ $=\tfrac{-[\nabla f(x^{+}(P,\beta))+\omega(x^{+}(P,\beta)-x)]\nabla f(x)^{\top}}{\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}},$ (49)
$\nabla_{\beta}h_{x,x^{-}}(P,\beta)$ $=\tfrac{\langle\nabla f(x^{+}(P,\beta))+\omega(x^{+}(P,\beta)-x),~x-x^{-}\rangle}{\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}}.$ (50)
Smoothness.
The difference of the gradients in $P$ and $\beta$ are respectively bounded by
$\|\nabla_{P}h_{x,x^{-}}(P_{1},\beta_{1})-\nabla_{P}h_{x,x^{-}}(P_{2},\beta_{2})\|_{F}={}$ $\tfrac{\|[\nabla f(x^{+}(P_{1},\beta_{1}))-\nabla f(x^{+}(P_{2},\beta_{2}))+\omega(x^{+}(P_{1},\beta_{1})-x^{+}(P_{2},\beta_{2}))]\nabla f(x)^{\top}\|_{F}}{\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}}$
$={}$ $\tfrac{\|\nabla f(x^{+}(P_{1},\beta_{1}))-\nabla f(x^{+}(P_{2},\beta_{2}))+\omega(x^{+}(P_{1},\beta_{1})-x^{+}(P_{2},\beta_{2}))\|\|\nabla f(x)\|}{\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}}$ (51)
$|\nabla_{\beta}h_{x,x^{-}}(P_{1},\beta_{1})-\nabla_{\beta}h_{x,x^{-}}(P_{2},\beta_{2})|={}$ $\tfrac{|\langle\nabla f(x^{+}(P_{1},\beta_{1}))-\nabla f(x^{+}(P_{2},\beta_{2}))+\omega(x^{+}(P_{1},\beta_{1})-x^{+}(P_{2},\beta_{2})),x-x^{-}\rangle|}{\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}}$
$\leq{}$ $\tfrac{\|\nabla f(x^{+}(P_{1},\beta_{1}))-\nabla f(x^{+}(P_{2},\beta_{2}))+\omega(x^{+}(P_{1},\beta_{1})-x^{+}(P_{2},\beta_{2}))\|\|x-x^{-}\|}{\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}},$ (52)
where (52) uses Cauchy’s inequality. Observe that the common norm in the numerator can be bounded by
$\|\nabla f(x^{+}(P_{1},\beta_{1}))-\nabla f(x^{+}(P_{2},\beta_{2}))+\omega(x^{+}(P_{1},\beta_{1})-x^{+}(P_{2},\beta_{2}))\|$
$\leq{}$ $\|\nabla f(x^{+}(P_{1},\beta_{1}))-\nabla f(x^{+}(P_{2},\beta_{2}))\|+\omega\|x^{+}(P_{1},\beta_{1})-x^{+}(P_{2},\beta_{2})\|$ (by $\|a+b\|\leq\|a\|+\|b\|$)
$\leq{}$ $(L+\omega)\|x^{+}(P_{1},\beta_{1})-x^{+}(P_{2},\beta_{2})\|$ (by $L$-smooth of $f$)
$={}$ $(L+\omega)\|(P_{2}-P_{1})\nabla f(x)+(\beta_{1}-\beta_{2})(x-x^{-})\|$
$\leq{}$ $(L+\omega)\|P_{1}-P_{2}\|\|\nabla f(x)\|+|\beta_{1}-\beta_{2}|\|x-x^{-}\|$ (by submultiplicativity)
$\leq{}$ $(L+\omega)\sqrt{\|P_{1}-P_{2}\|^{2}_{F}+\tfrac{1}{L^{2}}(\beta_{1}-\beta_{2})^{2}}\sqrt{\|\nabla f(x)\|^{2}+L^{2}\|x-x^{-}\|^{2}}$ (by Cauchy’s inequality)
$={}$ $(L+\omega)\sqrt{\|\nabla f(x)\|^{2}+L^{2}\|x-x^{-}\|^{2}}\|(P_{1},\beta_{1})-(P_{2},\beta_{2})\|.$ (by defintion of $\|(P,\beta)\|$ norm)
Then (51) and (52) can be further bounded by
$\|\nabla_{P}h_{x,x^{-}}(P_{1},\beta_{1})-\nabla_{P}h_{x,x^{-}}(P_{2},\beta_{2})\|_{F}$ $\leq(L+\omega)\tfrac{\sqrt{\|\nabla f(x)\|^{2}+L^{2}\|x-x^{-}\|^{2}}\|\nabla f(x)\|}{\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}}\|(P_{1},\beta_{1})-(P_{2},\beta_{2})\|,$
$|\nabla_{\beta}h_{x,x^{-}}(P_{1},\beta_{1})-\nabla_{\beta}h_{x,x^{-}}(P_{2},\beta_{2})|$ $\leq(L+\omega)\tfrac{\sqrt{\|\nabla f(x)\|^{2}+L^{2}\|x-x^{-}\|^{2}}\|x-x^{-}\|}{\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}}\|(P_{1},\beta_{1})-(P_{2},\beta_{2})\|.$
Hence the dual norm $\|\cdot\|_{\ast}$ of the gradient difference is bounded by
$\|(\nabla_{P}h_{x,x^{-}}(P_{1},\beta_{1})-\nabla_{P}h_{x,x^{-}}(P_{2},\beta_{2}),\nabla_{\beta}h_{x,x^{-}}(P_{1},\beta_{1})-\nabla_{\beta}h_{x,x^{-}}(P_{2},\beta_{2}))\|_{\ast}^{2}$
$={}$ $\|\nabla_{P}h_{x,x^{-}}(P_{1},\beta_{1})-\nabla_{P}h_{x,x^{-}}(P_{2},\beta_{2})\|^{2}_{F}+L^{2}|\nabla_{\beta}h_{x,x^{-}}(P_{1},\beta_{1})-\nabla_{\beta}h_{x,x^{-}}(P_{2},\beta_{2})|^{2}$
$\leq{}$ $(L+\omega)^{2}\tfrac{[\|\nabla f(x)\|^{2}+L^{2}\|x-x^{-}\|^{2}]^{2}}{[\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}]^{2}}\|(P_{1},\beta_{1})-(P_{2},\beta_{2})\|^{2}$
$\leq{}$ $(L+\omega)^{2}\max\{\tfrac{4L^{4}}{\tau^{2}},1\}\|(P_{1},\beta_{1})-(P_{2},\beta_{2})\|^{2}$ (by $\tfrac{a^{2}+\gamma b^{2}}{a^{2}+\delta b^{2}}\leq\max\{\tfrac{\gamma}{\delta},1\}$)
$={}$ $(L+\omega)^{2}\|(P_{1},\beta_{1})-(P_{2},\beta_{2})\|^{2},$ (by $\tau\geq 2L^{2}$)
which implies $(L+\omega)$-smoothness of $h_{x,x^{-}}(P,\beta)$. Finally, as a by-product relation that will be later used, we derive the dual norm $\|\cdot\|_{\ast}$ of the gradient is bounded by
$\|(\nabla_{P}h_{x,x^{-}}(P,\beta),\nabla_{\beta}h_{x,x^{-}}(P,\beta))\|^{2}_{\ast}$ $=\|\nabla_{P}h_{x,x^{-}}(P,\beta)\|^{2}+L^{2}\|\nabla_{\beta}h_{x,x^{-}}(P,\beta)\|^{2}$
$\leq\tfrac{\|\nabla f(x^{+}(P,\beta))+\omega(x^{+}(P,\beta)-x)\|^{2}[\|\nabla f(x)\|^{2}+L^{2}\|x-x^{-}\|^{2}]}{[\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}]^{2}}$ (by (49) and (50))
$\leq\max\{\tfrac{2L^{2}}{\tau},1\}\tfrac{\|\nabla f(x^{+}(P,\beta))+\omega(x^{+}(P,\beta)-x)\|^{2}}{\|\nabla f(x)\|^{2}+\frac{\tau}{2}\|x-x^{-}\|^{2}}.$ (53)
A.5 Proof of Theorem 2.2
Note that the initial state $z^{1}$ with $z_{2}^{1}=z_{1}^{1}$ implies that the initial potential is the initial suboptimality:
$\varphi_{\omega}(z^{1})=f(z^{1})-f^{\star}+\tfrac{\omega}{2}\|z_{1}^{1}-z_{2}^{1}\|^{2}=f(z^{1})-f^{\star}.$
Convex $f$.
Observe that
$\varphi_{\omega}(z^{K+1})=\frac{1}{\sum^{K}_{k=1}\big[\frac{\varphi_{\omega}(z^{k})-\varphi_{\omega}(z^{k+1})}{\varphi_{\omega}(z^{k+1})\varphi_{\omega}(z^{k})}\big]+\tfrac{1}{\varphi_{\omega}(z^{1})}}=\frac{1}{\sum^{K}_{k=1}-b_{k}\big[\frac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{\varphi_{\omega}(z^{k+1})\varphi_{\omega}(z^{k})}\big]+\tfrac{1}{f(z^{1})-f^{\star}}}.$ (54)
Note that $b_{k}\leq 0$ since the algorithm guarantees that the potential is monotone $\varphi_{\omega}(z^{k+1})\leq\varphi_{\omega}(z^{k})$. To further upper bound $\varphi_{\omega}(z^{K+1})$, we need to lower bound the ratio
$\tfrac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{\varphi_{\omega}(z^{k+1})\varphi_{\omega}(z^{k})}\geq\tfrac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{\varphi_{\omega}(z^{k})^{2}}=\tfrac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{\big[f(z^{k})-f^{\star}+\frac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}\big]^{2}}.$
Case 1. $f(z^{k})-f^{\star}\geq\tfrac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}$. Observe that $z_{1}^{k}\in\{x:f(x)\leq f(z^{1})\}$ for all $k\geq 1$ since the monotonicity of potential guarantees
$f(z^{k})-f^{\star}\leq f(z^{k})-f^{\star}+\tfrac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}=\varphi_{\omega}(z^{k})\leq\varphi_{\omega}(z^{1})=f(z^{1})-f^{\star}.$
Then $\tfrac{f(z^{k})-f^{\star}}{\|\nabla f(z^{k})\|}\leq\|z_{1}^{k}-x^{\star}\|\leq\Delta\coloneqq\max_{x\in\{x:f(x)\leq f(z^{1})\}}\min_{x^{\star}\in\mathcal{X}^{\star}}\|x-x^{\star}\|$ and hence
$\tfrac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{[f(z^{k})-f^{\star}+\frac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}]^{2}}\geq\tfrac{\|\nabla f(z^{k})\|^{2}}{4[f(z^{k})-f^{\star}]^{2}}\geq\tfrac{1}{4\Delta^{2}}.$ (55)
Case 2. $f(z^{k})-f^{\star}\leq\tfrac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}$. Observe that
$\tfrac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}\leq\varphi_{\omega}(z^{k})\leq\varphi_{\omega}(z^{1})=f(z^{1})-f^{\star}.$ (56)
Using (56) and the choice of $\tau\geq 2L\omega$, we have
$\tfrac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{[f(z^{k})-f^{\star}+\frac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}]^{2}}\geq\tfrac{\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{\omega^{2}\|z_{1}^{k}-z_{2}^{k}\|^{4}}=\tfrac{\tau}{2\omega^{2}}\tfrac{1}{\|z_{1}^{k}-z_{2}^{k}\|^{2}}\geq\tfrac{L}{2[f(z^{1})-f^{\star}]}.$ (57)
Using (54) and the two lower bounds (55) and (57), the suboptimality can be bounded by
$f(z^{K+1})-f^{\star}\leq\varphi_{\omega}(z^{K+1})\leq\tfrac{f(z^{1})-f^{\star}}{1+V\sum^{K}_{k=1}-b_{k}},$
where $V\coloneqq\min\big\{\tfrac{f(z^{1})-f^{\star}}{4\Delta^{2}},\tfrac{L}{2}\big\}$.
$\mu$-strongly convex $f$.
Observe that
$\tfrac{\varphi_{\omega}(z^{K+1})}{\varphi_{\omega}(z^{1})}=\textstyle\prod_{k=1}^{K}\tfrac{\varphi_{\omega}(z^{k+1})}{\varphi_{\omega}(z^{k})}$ $\leq\big(\tfrac{1}{K}\textstyle\sum_{k=1}^{K}\tfrac{\varphi_{\omega}(z^{k+1})}{\varphi_{\omega}(z^{k})}\big)^{K}$
$=(1+\tfrac{1}{K}\textstyle\sum_{k=1}^{K}\tfrac{\varphi_{\omega}(z^{k+1})-\varphi_{\omega}(z^{k})}{\varphi_{\omega}(z^{k})})^{K}$
$=(1-\tfrac{1}{K}\textstyle\sum_{k=1}^{K}-b_{k}\tfrac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{f(z^{k})-f^{\star}+\frac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}})^{K}.$ (58)
To further upper bound (58), we lower bound the ratio $\tfrac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{f(z^{k})-f^{\star}+\frac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}$ by case analysis.
Case 1. $f(z^{k})-f^{\star}\leq\tfrac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}$. Then
$\tfrac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{f(z^{k})-f^{\star}+\frac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}\geq\tfrac{\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{\omega\|z_{1}^{k}-z_{2}^{k}\|^{2}}=\tfrac{\tau}{2\omega}.$ (59)
Case 2. $f(z^{k})-f^{\star}\geq\tfrac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}$. Then using $\tfrac{\|\nabla f(z^{k})\|^{2}}{f(z_{1}^{k})-f^{\star}}\geq 2\mu$ to obtain
$\tfrac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{f(z^{k})-f^{\star}+\frac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}\geq\tfrac{\|\nabla f(z^{k})\|^{2}}{2[f(z^{k})-f^{\star}]}\geq\mu.$ (60)
Since $\omega$ and $\tau$ are chosen such that $\tfrac{\tau}{2\omega}\geq L\geq\mu$, two lower bounds (59)–(60) together imply
$\tfrac{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}{f(z^{k})-f^{\star}+\frac{\omega}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}\geq\min\{\tfrac{\tau}{2\omega},\mu\}=\mu,$
and hence we conclude from (58) that
$f(z^{K+1})-f^{\star}\leq\varphi_{\omega}(z^{K+1})\leq\varphi_{\omega}(z^{1})\big(1-\tfrac{\mu}{K}\textstyle\sum_{k=1}^{K}-b_{k}\big)^{K}=[f(z^{1})-f^{\star}]\big(1-\tfrac{\mu}{K}\textstyle\sum_{k=1}^{K}-b_{k}\big)^{K}.$
A.6 Proof of Theorem 2.3
Recall that the per iteration progress of OSGM-HB is
$b_{k}\coloneqq\tfrac{\varphi_{\omega}(z^{k+1})-\varphi_{\omega}(z^{k})}{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}.$
The scheduler proposes the state
$z^{k+1/2}\coloneqq(z^{k})^{+}=(z_{1}^{k}-P_{k}\nabla f(z^{k})+\beta_{k}(z_{1}^{k}-z_{2}^{k}),z_{1}^{k}).$
Monotone Lookahead OSGM-HB determines the next state $z^{k+1}$ such that
$\varphi_{\omega}(z^{k+1})\leq\min\{\varphi_{\omega}(\mathcal{M}_{\text{look}}(z^{k+1/2})),\varphi_{\omega}(z^{k})\},$ (61)
where $\mathcal{M}_{\text{look}}(z)\coloneqq(z_{1}-\tfrac{1}{L+\omega}\nabla_{1}\varphi_{\omega}(z),z_{2})$ The next lemma quantifies the improvement on iteration progress by this landscape action.
Lemma A.3.
Suppose $\tau\geq 2L^{2}$. Then the monotone lookahead landscape action (61) guarantees
$b_{k}\leq\min\{h_{z^{k}}(P_{k},\beta_{k})-\tfrac{1}{2(L+\omega)}\|\nabla h_{z^{k}}(P_{k},\beta_{k})\|_{\ast}^{2},0\}.$ (62)
Proof.
The monotone landscape action guarantees $\varphi_{\omega}(z^{k+1})\leq\min\{\varphi_{\omega}(\mathcal{M}_{\text{look}}(z^{k+1/2})),\varphi_{\omega}(z^{k})\},$ and hence
$b_{k}\coloneqq\tfrac{\varphi_{\omega}(z^{k+1})-\varphi_{\omega}(z^{k})}{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}\leq\min\{\tfrac{\varphi_{\omega}(\mathcal{M}_{\text{look}}(z^{k+1/2}))-\varphi_{\omega}(z^{k})}{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}},0\}.$ (63)
Descent lemma on $(L+\omega)$-smooth potential $\varphi_{\omega}(\cdot,z_{2}^{k+1/2})$ implies
$\varphi_{\omega}\big(\mathcal{M}_{\text{look}}(z^{k+1/2})\big)$ $\leq\varphi_{\omega}(z^{k+1/2})-\tfrac{1}{2(L+\omega)}\|\nabla_{1}\varphi_{\omega}(z^{k+1/2})\|^{2}$
$=\varphi(z^{k+1/2})-\tfrac{1}{2(L+\omega)}\|\nabla f(z^{k+1/2})+\omega(z_{1}^{k+1/2}-z_{2}^{k+1/2})\|^{2}.$
Subtracting both sides by $\varphi_{\omega}(z^{k})$ and dividing by $\|\nabla f(z^{k})\|^{2}+\tfrac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}$, we obtain
$\tfrac{\varphi_{\omega}(\mathcal{M}_{\text{look}}(z^{k+1/2}))-\varphi_{\omega}(z^{k})}{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}$ $\leq\tfrac{\varphi_{\omega}(z^{k+1/2})-\varphi_{\omega}(z^{k})}{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}-\tfrac{1}{2(L+\omega)}\tfrac{\|\nabla f(z^{k+1/2})+\omega(z_{1}^{k+1/2}-z_{2}^{k+1/2})\|^{2}}{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}$
$=h_{z^{k}}(P_{k},\beta_{k})-\tfrac{1}{2(L+\omega)}\|\nabla h_{z^{k}}(P_{k},\beta_{k})\|_{\ast}^{2},$ (64)
where (64) uses the inequality of gradient norm from (53) and that $\tau\geq 2L^{2}$:
$\|\nabla h_{z^{k}}(P_{k},\beta_{k})\|_{\ast}^{2}\leq\tfrac{\|\nabla f(z^{k+1/2})+\omega(z_{1}^{k+1/2}-z_{2}^{k+1/2})\|^{2}}{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}.$
Combining the two inequalities (63) and (64) completes the proof. ∎
To bound the cumulative feedback $\sum_{k=1}^{K}h_{z^{k}}(P_{k},\beta_{k})$, we use the analysis of online gradient descent.
Lemma A.4.
Consider online gradient descent on heavy-ball feedback
$P_{k+1}$ $=\Pi_{\mathcal{P}}[P_{k}-\eta_{P}\nabla h_{z^{k}}(P_{k},\beta_{k})],$
$\beta_{k+1}$ $=\Pi_{\mathcal{B}}[\beta_{k}-\eta_{\beta}\nabla h_{z^{k}}(P_{k},\beta_{k})],$
with $\eta_{P}=\eta$ and $\eta_{\beta}=L^{2}\eta$ for some $\eta>0$. Then for any $(\hat{P},\hat{\beta})\in\mathcal{P}\times\mathcal{B}$, the sequence $\{(P_{k},\beta_{k})\}$ satisfies
$\sum_{k=1}^{K}h_{z^{k}}(P_{k},\beta_{k})\leq\sum_{k=1}^{K}h_{z^{k}}(\hat{P},\hat{\beta})+\tfrac{1}{2\eta}\|(P_{1}-\hat{P},\beta_{1}-\hat{\beta})\|^{2}+\tfrac{\eta}{2}\sum_{k=1}^{K}\|\nabla h_{z^{k}}(P_{k},\beta_{k})\|^{2}_{\ast}.$
Proof.
The proof is similar to the analysis of online gradient descent under Frobenius norm. Note that our algorithm finally takes $\mathcal{P}=\mathbb{R}^{n\times n}$, and the analysis here is more general. Observe that
$\|(P_{k+1}-\hat{P},\beta_{k+1}-\hat{\beta})\|^{2}$ $=\|\Pi_{\mathcal{P}}[P_{k}-\eta_{P}\nabla_{P}h_{z^{k}}(P_{k},\beta_{k})]-\hat{P}\|^{2}_{F}+\tfrac{1}{L^{2}}|\Pi_{\mathcal{B}}[\beta_{k}-\eta_{\beta}\nabla_{\beta}h_{z^{k}}(P_{k},\beta_{k})]-\hat{\beta}|^{2}$
$\leq\|P_{k}-\hat{P}-\eta_{P}\nabla_{P}h_{z^{k}}(P_{k},\beta_{k})\|^{2}_{F}+\tfrac{1}{L^{2}}|\beta_{k}-\hat{\beta}-\eta_{\beta}\nabla_{\beta}h_{z^{k}}(P_{k},\beta_{k})|^{2}$
$=\|P_{k}-\hat{P}\|^{2}_{F}-2\eta_{P}\langle P_{k}-\hat{P},\nabla_{P}h_{z^{k}}(P_{k},\beta_{k})\rangle+\eta^{2}_{P}\|\nabla_{P}h_{z^{k}}(P_{k},\beta_{k})\|^{2}$
$\quad+\tfrac{1}{L^{2}}|\beta_{k}-\hat{\beta}|^{2}-2\tfrac{\eta_{\beta}}{L^{2}}\langle\beta_{k}-\hat{\beta},\nabla_{\beta}h_{z^{k}}(P_{k},\beta_{k})\rangle+\tfrac{\eta^{2}_{\beta}}{L^{2}}|\nabla_{\beta}h_{z^{k}}(P_{k},\beta_{k})|^{2}$
$=\|(P_{k}-\hat{P},\beta_{k}-\hat{\beta})\|^{2}+\eta^{2}[\|\nabla_{P}h_{z^{k}}(P_{k},\beta_{k})\|^{2}+L^{2}|\nabla_{\beta}h_{z^{k}}(P_{k},\beta_{k})|^{2}]$
$\quad-2\eta[\langle P_{k}-\hat{P},\nabla_{P}h_{z^{k}}(P_{k},\beta_{k})\rangle+\langle\beta_{k}-\hat{\beta},\nabla_{\beta}h_{z^{k}}(P_{k},\beta_{k})\rangle]$
[…]
which rearranges to
$h_{z^{k}}(P_{k},\beta_{k})\leq h_{z^{k}}(\hat{P},\hat{\beta})+\tfrac{1}{2\eta}[\|(P_{k}-\hat{P},\beta_{k}-\hat{\beta})\|^{2}-\|(P_{k+1}-\hat{P},\beta_{k+1}-\hat{\beta})\|^{2}]+\tfrac{\eta}{2}\|\nabla h_{z^{k}}(P_{k},\beta_{k})\|_{\ast}^{2}.$
Telescope from $k=1,\ldots,K$ to obtain
$\textstyle\sum_{k=1}^{K}h_{z^{k}}(P_{k},\beta_{k})$ $\leq\textstyle\sum_{k=1}^{K}h_{z^{k}}(\hat{P},\hat{\beta})+\tfrac{1}{2\eta}\|(P_{1}-\hat{P},\beta_{1}-\hat{\beta})\|^{2}+\tfrac{\eta}{2}\textstyle\sum_{k=1}^{K}\|\nabla h_{z^{k}}(P_{k},\beta_{k})\|_{\ast}^{2}.$
Proof of Theorem˜2.3.
Note that our choice of $\tau=16L^{2}$ satisfies the condition $\tau\geq 2L^{2}$ in Lemma˜A.3. By Lemma˜A.3 and Lemma˜A.4, the cumulative progress of Monotone Lookahead OSGM-HB with $\eta\leq\tfrac{1}{L+\omega}$ can be bounded by
$\textstyle\sum_{k=1}^{K}b_{k}$ $\leq\textstyle\sum_{k=1}^{K}\min\{h_{z^{k}}(P_{k},\beta_{k})-\tfrac{1}{2(L+\omega)}\|\nabla h_{z^{k}}(P_{k},\beta_{k})\|_{\ast}^{2},0\}$ (by Lemma A.3)
$\leq\min\{\textstyle\sum_{k=1}^{K}[h_{z^{k}}(P_{k},\beta_{k})-\tfrac{\eta}{2}\|\nabla h_{z^{k}}(P_{k},\beta_{k})\|_{\ast}^{2}],0\}$ (by $\eta\leq\tfrac{\tau}{2L^{2}(L+\omega)}$)
$\leq\min\{\textstyle\sum_{k=1}^{K}h_{z^{k}}(\hat{P},\hat{\beta})+\tfrac{1}{2\eta}\|(P_{1}-\hat{P},\beta_{1}-\hat{\beta})\|^{2},0\}.$ (by Lemma A.4)
Choose the online gradient stepsize $\eta=\tfrac{1}{L+\omega}$ and denote
$\rho_{K}(\hat{P},\hat{\beta})\coloneqq\tfrac{1}{2\eta}\|(P_{1}-\hat{P},\beta_{1}-\hat{\beta})\|^{2}$
Then the bound on average progress of Monotone Lookahead OSGM-HB becomes
$\tfrac{1}{K}\textstyle\sum_{k=1}^{K}b_{k}\leq\min\{\textstyle\sum_{k=1}^{K}h_{z^{k}}(\hat{P},\hat{\beta})+\tfrac{\rho_{K}(\hat{P},\hat{\beta})}{K},0\}.$ (65)
Our choice of parameters $\omega=3L$ and $\tau=16L^{2}$ also satisfies the condition $\tau\geq 2L\omega$ in Theorem˜2.2. Plug in the bound (65) into the heavy-ball reductions (9)–(10) (Theorem˜2.2) to obtain
$\text{(convex)}\quad f(z^{K+1})-f^{\star}$ $\leq\tfrac{f(z^{1})-f^{\star}}{KV\max\big\{\frac{1}{K}\sum_{k=1}^{K}-h_{z^{k}}(\hat{P},\hat{\beta})-\frac{\rho_{K}(\hat{P},\hat{\beta})}{K},0\big\}+1},$ (66)
$(\mu\text{-strongly convex)}\quad f(z^{K+1})-f^{\star}$ $\leq[f(z^{1})-f^{\star}]\big(1-\mu\max\big\{\tfrac{1}{K}\textstyle\sum_{k=1}^{K}-h_{z^{k}}(\hat{P},\hat{\beta})-\frac{\rho_{K}(\hat{P},\hat{\beta})}{K},0\big\}\big)^{K}.$ (67)
Lemma˜2.3 with $\alpha=\tfrac{1}{4L}$ and $\beta=\tfrac{1}{2}$ implies $\omega=3L$ and guarantees
$\varphi_{w}\big((z^{k})^{+}(\tfrac{1}{4L}I,\tfrac{1}{2})\big)$ $=\varphi_{w}\big(z_{1}^{k}-\tfrac{1}{4L}\nabla f(z^{k})+\tfrac{1}{2}(z_{1}^{k}-z_{2}^{k}),z_{1}^{k}\big)$
$\leq\varphi_{w}(z^{k})-\tfrac{1}{8L}[\|\nabla f(z^{k})\|^{2}+8L^{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}]$
$=\varphi_{w}(z^{k})-\tfrac{1}{8L}[\|\nabla f(z^{k})\|^{2}+\tfrac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}],$ (by $\tau=16L^{2}$)
which rearranges to
$h_{z^{k}}(\tfrac{1}{4L}I,\tfrac{1}{2})=\tfrac{\varphi_{\omega}((z^{k})^{+}(\frac{1}{4L}I,\frac{1}{2}))-\varphi_{\omega}(z^{k})}{\|\nabla f(z^{k})\|^{2}+\frac{\tau}{2}\|z_{1}^{k}-z_{2}^{k}\|^{2}}\leq-\tfrac{1}{8L}\text{, for all $k$.}$ (68)
Plug in $(\hat{P},\hat{\beta})=(\tfrac{1}{4L}I,\tfrac{1}{2})$ into (66) and (67) and further bound by (68) completes the proof.
Appendix B Proof of results in Section 3
B.1 Proof of Lemma 3.1
Since $h_{x}(P)$ is $L$-smooth, it is at most $L$-weakly convex. The second result needs several auxiliary lemmas.
Lemma B.1 (Hessian of PL-functions).
Suppose $f$ is $L$-smooth, has $H$-Lipschitz Hessian, and is $\mu$-PL, then $\lambda_{\min}(\nabla^{2}f(x))\geq-\mu^{-1}H\|\nabla f(x)\|$.
Proof.
Let $x^{\star}=\Pi_{\mathcal{X}^{\star}}[x]$. Using $H$-Lipschitz Hessian, we have
$\nabla^{2}f(x)={}$ $\nabla^{2}f(x^{\star})+\nabla^{2}f(x)-\nabla^{2}f(x^{\star})$
$\succeq{}$ $\nabla^{2}f(x^{\star})-\|\nabla^{2}f(x)-\nabla^{2}f(x^{\star})\|\cdot I$
$\succeq{}$ $\nabla^{2}f(x^{\star})-H\|x-x^{\star}\|\cdot I.$ ($H$-Lipschitz Hessian)
Since PL condition implies quadratic growth [21], $\tfrac{1}{2\mu}\|\nabla f(x)\|^{2}\geq f(x)-f^{\star}\geq\tfrac{\mu}{2}\|x-x^{\star}\|^{2}$ and $\|x-x^{\star}\|\leq\tfrac{1}{\mu}\|\nabla f(x)\|$. Hence $\nabla^{2}f(x)\succeq\nabla^{2}f(x^{\star})-\tfrac{H}{\mu}\|\nabla f(x)\|\cdot I\succeq-\mu^{-1}H\|\nabla f(x)\|\cdot I$ and this completes the proof. ∎
Lemma B.2 (Hessian of $h_{x}(P)$).
Suppose $f$ has $H$-Lipschitz Hessian and $x\not\in\mathcal{S}^{\star}$, then
$\|\nabla^{2}h_{x}(P_{1})-\nabla^{2}h_{x}(P_{2})\|\leq H\|\nabla f(x)\|\|P_{1}-P_{2}\|.$
Proof.
Recall that $\nabla h_{x}(P)=\tfrac{\nabla f(x-P\nabla f(x))\nabla f(x)^{\top}}{\|\nabla f(x)\|^{2}}$ and $\nabla^{2}h_{x}(P)=(gg^{\top})\otimes\nabla^{2}f(x-P\nabla f(x))$ where $g=\tfrac{\nabla f(x)}{\|\nabla f(x)\|}$. Since $\|g\|=1$, $H$-Lipschitz Hessian implies
$\|\nabla^{2}h_{x}(P_{1})-\nabla^{2}h_{x}(P_{2})\|\leq\|\nabla^{2}f(x-P_{1}\nabla f(x))-\nabla^{2}f(x-P_{2}\nabla f(x))\|\\ \leq H\|\nabla f(x)\|\|P_{1}-P_{2}\|$
and this completes the proof. ∎
Proof of Lemma˜3.1.
Take any fixed $\bar{P}\in\mathcal{P}$. Using Lemma˜B.2, we deduce that
$\nabla^{2}h_{x}(P)\succeq{}$ $\nabla^{2}h_{x}(\bar{P})-H\|\nabla f(x)\|\|P-\bar{P}\|\cdot I$
$\succeq{}$ $[\lambda_{\min}(\nabla^{2}f(x-\bar{P}\nabla f(x)))-H\|\nabla f(x)\|\|P-\bar{P}\|]\cdot I$
$\succeq{}$ $[\lambda_{\min}(\nabla^{2}f(x-\bar{P}\nabla f(x)))-H\|\nabla f(x)\|D]\cdot I$
When $f$ is additionally $\mu$-PL and $\bar{P}=0$, we can further apply $\lambda_{\min}(\nabla^{2}f(x-\bar{P}\nabla f(x)))\geq-\mu^{-1}H\|\nabla f(x)\|$ from Lemma˜B.1 and this completes the proof.
B.2 Proof of Lemma 3.2
The expression of gradient follows by definition $\nabla h_{x}^{\lambda}(P)=\nabla h_{x}(P)+\lambda(P-\tfrac{1}{L}I)$. $h_{x}^{\lambda}(\tfrac{1}{L}I)=h_{x}(\tfrac{1}{L}I)\leq-\tfrac{1}{2L}$ follows from the descent lemma. Using Lemma˜3.1, local convexity of $h_{x}^{\lambda}(P)$ follows immediately and the smoothness constant follows by adding $L$, the smoothness of $h_{x}(P)$, and $\lambda$.
B.3 Proof of Theorem 3.1
These reductions have appeared in [13], and we repeat the proofs for completeness. The first relation (15) follows by definition:
$(\textstyle\sum_{k=1}^{K}-h_{k})\cdot\displaystyle\min_{1\leq k\leq K}\|\nabla f(x^{k})\|^{2}\leq{}$ $\textstyle\sum_{k=1}^{K}-h_{k}\|\nabla f(x^{k})\|^{2}$ (by $h_{k}\geq 0$)
$={}$ $-[\textstyle\sum_{k=1}^{K}f(x^{k+1})-f(x^{k})]$ (by definition of $h_{k}$)
$={}$ $f(x^{1})-f(x^{K+1})$
$\leq{}$ $f(x^{1})-f^{\star}.$ (by $f(x^{K+1})\leq f^{\star}$)
and the second relation (16) follows from:
$\tfrac{f(x^{K+1})-f^{\star}}{f(x^{1})-f^{\star}}\leq{}$ $(\tfrac{1}{K}\textstyle\sum_{k=1}^{K}\tfrac{f(x^{k+1})-f^{\star}}{f(x^{k})-f^{\star}})^{K}$ (by AM-GM inequality)
$={}$ $(1+\tfrac{1}{K}\textstyle\sum_{k=1}^{K}\tfrac{f(x^{k+1})-f(x^{k})}{f(x^{k})-f^{\star}})^{K}$
$={}$ $(1+\tfrac{1}{K}\textstyle\sum_{k=1}^{K}\tfrac{h_{k}\|\nabla f(x^{k})\|^{2}}{f(x^{k})-f^{\star}})^{K}$ (by definition of $h_{k}$)
$\leq{}$ $(1-\tfrac{2\mu}{K}\textstyle\sum_{k=1}^{K}-h_{k})^{K}$ (by $h_{k}\leq 0$ and $\frac{\|\nabla f(x^{k})\|^{2}}{f(x^{k})-f^{\star}}\geq 2\mu$)
and this completes the proof.
B.4 Proof of Lemma 3.3
Recall that $x^{k+1}=\mathcal{M}(x^{k+1/2},x^{k})$ and $x^{k+1/2}=x^{k}-P_{k}\nabla f(x^{k})$. The per iteration progress and (regularized) feedback are
$h_{k}=\tfrac{f(x^{k+1})-f(x^{k})}{\|\nabla f(x^{k})\|^{2}},\quad h_{x^{k}}(P_{k})=\tfrac{f(x^{k+1/2})-f(x^{k})}{\|\nabla f(x^{k})\|^{2}},\quad h_{x^{k}}^{\lambda}(P_{k})=h_{x^{k}}(P_{k})+\tfrac{\lambda}{2}\|P_{k}-\tfrac{1}{L}I\|_{F}^{2}.$
  • Monotone landscape satisfies $f(x^{k+1})\leq\min\{f(x^{k+1/2}),f(x^{k})\}$ and thus
    $h_{k}=\tfrac{f(x^{k+1})-f(x^{k})}{\|\nabla f(x^{k})\|^{2}}\leq\min\{h_{x^{k}}(P_{k}),0\}=\min\{h_{x^{k}}^{\lambda}(P_{k})-\tfrac{\lambda}{2}\|P_{k}-\tfrac{1}{L}I\|_{F}^{2},0\}.$
  • Monotone lookahead landscape satisfies $f(x^{k+1})\leq\min\{f(x^{k+1/2}-\tfrac{1}{L}\nabla f(x^{k+1/2})),0\}$ and by descent lemma, $f(x^{k+1/2}-\tfrac{1}{L}\nabla f(x^{k+1/2}))\leq f(x^{k+1/2})-\tfrac{1}{2L}\|\nabla f(x^{k+1/2})\|^{2}$ and
    $h_{k}={}$ $\tfrac{f(x^{k+1})-f(x^{k})}{\|\nabla f(x^{k})\|^{2}}$
    $\leq{}$ $\tfrac{f(x^{k+1/2})-f(x^{k})}{\|\nabla f(x^{k})\|^{2}}-\tfrac{1}{2L}\tfrac{\|\nabla f(x^{k+1/2})\|^{2}}{\|\nabla f(x^{k})\|^{2}}$ (by descent lemma)
    $={}$ $h_{x^{k}}(P_{k})-\tfrac{1}{2L}\|\nabla h_{x^{k}}(P_{k})\|_{F}^{2}$ (by $\|\nabla h_{x^{k}}(P_{k})\|_{F}^{2}=\tfrac{\|\nabla f(x^{k+1/2})\|^{2}}{\|\nabla f(x^{k})\|^{2}}$)
    $={}$ $h_{x^{k}}^{\lambda}(P_{k})-\tfrac{\lambda}{2}\|P_{k}-\tfrac{1}{L}I\|_{F}^{2}-\tfrac{1}{2L}\|\nabla h_{x^{k}}(P_{k})\|_{F}^{2}$
    $={}$ $h_{x^{k}}^{\lambda}(P_{k})-\tfrac{1}{4L}\|\nabla h_{x^{k}}^{\lambda}(P_{k})\|_{F}^{2}+\tfrac{1}{4L}[\|\nabla h_{x^{k}}^{\lambda}(P_{k})\|_{F}^{2}-2\lambda L\|P_{k}-\tfrac{1}{L}I\|_{F}^{2}-2\|\nabla h_{x^{k}}(P_{k})\|_{F}^{2}].$
    Notice that when $\lambda\leq L$,
    $\|\nabla h_{x^{k}}^{\lambda}(P_{k})\|_{F}^{2}={}$ $\|\nabla h_{x^{k}}(P_{k})+\lambda(P_{k}-\tfrac{1}{L}I)\|_{F}^{2}$
    $\leq{}$ $2\|\nabla h_{x^{k}}(P_{k})\|_{F}^{2}+2\lambda^{2}\|P_{k}-\tfrac{1}{L}I\|_{F}^{2}$ (by $\|P+Q\|_{F}^{2}\leq 2\|P\|_{F}^{2}+2\|Q\|_{F}^{2}$)
    $\leq{}$ $2\|\nabla h_{x^{k}}(P_{k})\|_{F}^{2}+2\lambda L\|P_{k}-\tfrac{1}{L}I\|_{F}^{2}$ (by $\lambda\leq L$)
    which implies $h_{k}\leq h_{x^{k}}^{\lambda}(P_{k})-\tfrac{1}{4L}\|\nabla h_{x^{k}}^{\lambda}(P_{k})\|_{F}^{2}$ and completes the proof.
B.5 Proof of Theorem 3.2
Recall the regret guarantee of online gradient descent: for any $\hat{P}\in\mathcal{P}$, we successively deduce that
$\|P_{k+1}-\hat{P}\|_{F}^{2}={}$ $\|\Pi_{\mathcal{P}}[P_{k}-\eta\nabla h_{x^{k}}^{\lambda_{k}}(P_{k})]-\hat{P}\|_{F}^{2}$
$\leq{}$ $\|P_{k}-\eta\nabla h_{x^{k}}^{\lambda_{k}}(P_{k})-\hat{P}\|_{F}^{2}$ (by non-expansiveness)
$={}$ $\|P_{k}-\hat{P}\|_{F}^{2}-2\eta\langle\nabla h_{x^{k}}^{\lambda_{k}}(P_{k}),P_{k}-\hat{P}\rangle+\eta^{2}\|\nabla h_{x^{k}}^{\lambda_{k}}(P_{k})\|^{2}_{F}$
$\leq{}$ $\|P_{k}-\hat{P}\|_{F}^{2}-2\eta[h_{x^{k}}^{\lambda_{k}}(P_{k})-h_{x^{k}}^{\lambda_{k}}(\hat{P})]+\eta^{2}\|\nabla h_{x^{k}}^{\lambda_{k}}(P_{k})\|^{2}_{F},$
where the last inequality holds since $P_{k},\hat{P}\in\mathcal{P}$ and our choice of $\{\lambda_{k}\}$ ensures $h_{x^{k}}^{\lambda_{k}}(P)$ is locally convex. Next, we rearrange the terms and obtain
$h_{x^{k}}^{\lambda_{k}}(P_{k})-h_{x^{k}}^{\lambda_{k}}(\hat{P})\leq\tfrac{1}{2\eta}[\|P_{k}-\hat{P}\|_{F}^{2}-\|P_{k+1}-\hat{P}\|_{F}^{2}]+\tfrac{\eta}{2}\|\nabla h_{x^{k}}^{\lambda_{k}}(P_{k})\|^{2}_{F}.$ (69)
Back to the algorithm analysis. First, according to Theorem˜3.1, the convergence of Monotone Lookahead OSGM-H follows from bounding the cumulative progress
$\min_{1\leq k\leq K}\|\nabla f(x^{k})\|^{2}\leq{}$ $\tfrac{f(x^{1})-f^{\star}}{K}\tfrac{1}{\frac{1}{K}\sum_{k=1}^{K}-h_{k}},$ (70)
$f(x^{K+1})-f^{\star}\leq{}$ $[f(x^{1})-f^{\star}](1-\tfrac{2\mu}{K}\textstyle\sum_{k=1}^{K}-h_{k})^{K}.$ (71)
Second, the cumulative progress $\sum_{k=1}^{K}h_{k}$ under monotone lookahead action is bounded by Lemma˜3.3
$\tfrac{1}{K}\textstyle\sum_{k=1}^{K}h_{k}\leq\tfrac{1}{K}\textstyle\sum_{k=1}^{K}\min\{h^{\lambda_{k}}_{x^{k}}(P_{k})-\tfrac{1}{4L}\|\nabla h^{\lambda_{k}}_{x^{k}}(P_{k})\|_{F}^{2},0\}\leq\min\{\tfrac{1}{K}\textstyle\sum_{k=1}^{K}h^{\lambda_{k}}_{x^{k}}(P_{k})-\tfrac{1}{4L}\|\nabla h^{\lambda_{k}}_{x^{k}}(P_{k})\|_{F}^{2},0\}.$ (72)
Third, the regret guarantee of online gradient descent (69) bounds the cumulative feedback $\sum_{k=1}^{K}h^{\lambda_{k}}_{x^{k}}(P_{k})$ by
$\sum_{k=1}^{K}h^{\lambda_{k}}_{x^{k}}(P_{k})\leq\textstyle\sum_{k=1}^{K}h_{x^{k}}^{\lambda_{k}}(\hat{P})+\tfrac{1}{2\eta}\|P_{1}-\hat{P}\|_{F}^{2}+\tfrac{\eta}{2}\textstyle\sum_{k=1}^{K}\|\nabla h_{x^{k}}^{\lambda_{k}}(P_{k})\|^{2}_{F}.$ (73)
Putting (72) and (73) together, the cumulative progress is bounded by
$\tfrac{1}{K}\textstyle\sum_{k=1}^{K}h_{k}\leq\min\{\textstyle\sum_{k=1}^{K}h_{x^{k}}^{\lambda_{k}}(\hat{P})+\tfrac{1}{2\eta}\|P_{1}-\hat{P}\|_{F}^{2}+(\tfrac{\eta}{2}-\tfrac{1}{4L})\textstyle\sum_{k=1}^{K}\|\nabla h_{x^{k}}^{\lambda_{k}}(P_{k})\|^{2}_{F},0\}.$
When $\eta\leq\tfrac{1}{2L}$, the gradient norm term vanishes. With $\eta=\tfrac{1}{2L}$ the relation simplifies to
$\tfrac{1}{K}\textstyle\sum_{k=1}^{K}h_{k}\leq\min\{\textstyle\sum_{k=1}^{K}h_{x^{k}}^{\lambda_{k}}(\hat{P})+L\|P_{1}-\hat{P}\|_{F}^{2},0\}=\min\{\textstyle\sum_{k=1}^{K}h_{x^{k}}(\hat{P})+\tfrac{\lambda}{2}\|P-\tfrac{1}{L}\|_{F}^{2}+L\|P_{1}-\hat{P}\|_{F}^{2},0\}.$
Plugging the relation back into (70) and (71) completes the proof. Also notice that when $P_{1}=\hat{P}=\tfrac{1}{L}I$, $h_{x}^{\lambda}(\hat{P})\leq-\tfrac{1}{2L}$ and $\|P_{1}-\hat{P}\|_{F}^{2}=0$.
Appendix C Proof of results in Section 5
C.1 Details of Subsection 5.1
In this section, we replace the online gradient descent step in Monotone Lookahead OSGM-H with online proximal point. For the generality of the result, we still use a matrix stepsize $P$.
1input initial point $x^{1}$, initial stepsize $P_{1}\in\mathbb{R}^{n\times n}$, online proximal regularizer $\eta>0$
2for $k=1,2,\dots$ do
3 $x^{k+1/2}=x^{k}-P_{k}\nabla f(x^{k})$
4   Choose $x^{k+1}$ that satisfies $f(x^{k+1})\leq\min\{f(x^{k+1/2}-\frac{1}{L}\nabla f(x^{k+1/2})),f(x^{k})\}$
5 $P_{k+1}=\operatornamewithlimits{arg\,min}_{P}\{h_{x^{k}}(P)+\tfrac{1}{2\eta}\|P-P_{k}\|_{F}^{2}\}$
6 end for
Algorithm 7 Monotone Lookahead OSGM-H with online proximal point update
To show the convergence of Algorithm˜7, we again proceed in three steps: 1) bounding suboptimality with progress 2) bounding progress with feedback 3) bounding feedback with online learning. Since we essentially use the same hypergradient feedback, step 1) and 2) are the same as in Part I. It suffices to prove the regret guarantee of online proximal point and combine it with the lookahead action. Recall the definition of progress $h_{k}=\tfrac{f(x^{k+1})-f(x^{k})}{\|\nabla f(x^{k})\|^{2}}$ and feedback $h_{x^{k}}(P_{k})=\tfrac{f(x^{k+1/2})-f(x^{k})}{\|\nabla f(x^{k})\|^{2}}$. We have, from [14, Lemma 4.1], that
Lemma C.1 (Feedback and progress [14]).
Let $f$ be $L$-smooth and convex. Then $h_{k}\leq\min\{h_{x^{k}}(P_{k})-\tfrac{1}{2L}\|\nabla h_{x^{k}}(P_{k})\|_{F}^{2},0\}$ with monotone lookahead action.
The proof requires an additional technical lemma that shows a relation between the value of the feedback function $h_{x}(P)$ and its Moreau envelope function $h_{x}^{1/\rho}(P)\coloneqq\min_{Q}\{h_{x}(Q)+\tfrac{\rho}{2}\|Q-P\|_{F}^{2}\}$.
Lemma C.2.
Let $h^{1/\rho}(x)\coloneqq\min_{z}\{h(z)+\tfrac{\rho}{2}\|z-x\|^{2}\}$ be the Moreau envelope of a differentiable convex function $f$. Then $h(x)-h^{1/\rho}(x)-\tfrac{1}{2\rho}\|\nabla h(x)\|^{2}\leq 0$.
Proof.
By convexity, we have, for all $z$, that
$h(z)+\tfrac{\rho}{2}\|z-x\|^{2}\geq{}$ $h(x)+\langle\nabla h(x),z-x\rangle+\tfrac{\rho}{2}\|z-x\|^{2}$
$\geq{}$ $\min_{d}\{h(x)+\langle\nabla h(x),d\rangle+\tfrac{\rho}{2}\|d\|^{2}\}$
$={}$ $h(x)-\tfrac{1}{2\rho}\|\nabla h(x)\|^{2}.$
Minimizing the LHS, $h^{1/\rho}(x)=\min_{z}\{h(z)+\tfrac{\rho}{2}\|z-x\|^{2}\}\geq h(x)-\tfrac{1}{2\rho}\|\nabla h(x)\|^{2}$. This completes the proof. ∎
Using the optimality condition on the proximal subproblem on Algorithm˜7 of Algorithm˜7, for any $\hat{P}$, we deduce that
$h_{x^{k}}(P_{k+1})+\tfrac{1}{2\eta}\|P_{k+1}-P_{k}\|_{F}^{2}\leq{}$ $h_{x^{k}}(\hat{P})+\tfrac{1}{2\eta}\|\hat{P}-P_{k}\|_{F}^{2}-\tfrac{1}{2\eta}\|P_{k+1}-\hat{P}\|_{F}^{2}.$
Re-arranging the terms, $h_{x^{k}}(P_{k+1})-h_{x^{k}}(\hat{P})\leq\tfrac{1}{2\eta}\|\hat{P}-P_{k}\|_{F}^{2}-\tfrac{1}{2\eta}\|P_{k+1}-P_{k}\|_{F}^{2}-\tfrac{1}{2\eta}\|P_{k+1}-\hat{P}\|_{F}^{2}$ and
[…]
$\leq{}$ $h_{x^{k}}(P_{k})-h_{x^{k}}(P_{k+1})+\tfrac{1}{2\eta}\|\hat{P}-P_{k}\|_{F}^{2}-\tfrac{1}{2\eta}\|P_{k+1}-P_{k}\|_{F}^{2}-\tfrac{1}{2\eta}\|P_{k+1}-\hat{P}\|_{F}^{2}$
$={}$ $h_{x^{k}}(P_{k})-[h_{x^{k}}(P_{k+1})+\tfrac{1}{2\eta}\|P_{k+1}-P_{k}\|_{F}^{2}]+\tfrac{1}{2\eta}[\|\hat{P}-P_{k}\|_{F}^{2}-\|P_{k+1}-\hat{P}\|_{F}^{2}]$
$={}$ $h_{x^{k}}(P_{k})-h_{x^{k}}^{\eta}(P_{k})+\tfrac{1}{2\eta}[\|\hat{P}-P_{k}\|_{F}^{2}-\|P_{k+1}-\hat{P}\|_{F}^{2}]$
$\leq{}$ $\tfrac{\eta}{2}\|\nabla h_{x^{k}}(P_{k})\|^{2}+\tfrac{1}{2\eta}[\|\hat{P}-P_{k}\|_{F}^{2}-\|P_{k+1}-\hat{P}\|_{F}^{2}]$
Finally, we bound the progress using the monotone lookahead action using Lemma˜C.1:
$\textstyle\sum_{k=1}^{K}h_{k}\leq$ $\textstyle\min\{\sum_{k=1}^{K}h_{x^{k}}(P_{k})-\tfrac{1}{2L}\|\nabla h_{x^{k}}(P_{k})\|_{F}^{2},0\}$
$\leq$ $\textstyle\min\{\sum_{k=1}^{K}h_{x^{k}}(\hat{P})+(\tfrac{\eta}{2}-\tfrac{1}{2L})\|\nabla h_{x^{k}}(P_{k})\|^{2}+\tfrac{1}{2\eta}\|P_{1}-\hat{P}\|_{F}^{2},0\}$
and for $\eta=\tfrac{1}{L}$, $\sum_{k=1}^{K}h_{k}\leq\sum_{k=1}^{K}h_{x^{k}}(\hat{P})+\tfrac{L}{2}\|P_{1}-\hat{P}\|_{F}^{2}$. Plugging the progress relation back into the reduction in Part I [14, Theorem 4.2] establishes the convergence.
C.2 Proof of Theorem 5.1
By definition of the AGD update, we deduce that
$y^{k}={}$ $x^{k}+\tfrac{1}{\sqrt{\kappa}+1}(z^{k}-x^{k})$
$={}$ $\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}x^{k}+\tfrac{1}{\sqrt{\kappa}+1}z^{k}$
$x^{k+1}-x^{\star}={}$ $y^{k}-\tfrac{1}{L}\nabla f(y^{k})-x^{\star}$
$={}$ $y^{k}-\tfrac{1}{L}(Ay^{k}-Ax^{\star})-x^{\star}$
$={}$ $(I-\tfrac{1}{L}A)y^{k}+\tfrac{1}{L}Ax^{\star}-x^{\star}$
$={}$ $(I-\tfrac{1}{L}A)(\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}x^{k}+\tfrac{1}{\sqrt{\kappa}+1}z^{k})+\tfrac{1}{L}Ax^{\star}-x^{\star}$
$={}$ $(I-\tfrac{1}{L}A)\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}(x^{k}-x^{\star})-(I-\tfrac{1}{L}A)\tfrac{1}{\sqrt{\kappa}+1}(z^{k}-x^{\star})$
$z^{k+1}-x^{\star}={}$ $(1-\tfrac{1}{\sqrt{\kappa}})(z^{k}-x^{\star})+\tfrac{1}{\sqrt{\kappa}}(y^{k}-\tfrac{1}{\mu}(Ay^{k}-Ax^{\star})-x^{\star})$
$={}$ $(1-\tfrac{1}{\sqrt{\kappa}})(z^{k}-x^{\star})+\tfrac{1}{\sqrt{\kappa}}(I-\tfrac{1}{\mu}A)(y^{k}-x^{\star})$
$={}$ $(1-\tfrac{1}{\sqrt{\kappa}})(z^{k}-x^{\star})+\tfrac{1}{\sqrt{\kappa}}(I-\tfrac{1}{\mu}A)(\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}x^{k}+\tfrac{1}{\sqrt{\kappa}+1}z^{k}-x^{\star})$
$={}$ $\tfrac{1}{\sqrt{\kappa}+1}(I-\tfrac{1}{\mu}A)(x^{k}-x^{\star})+\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}(I-\tfrac{1}{L}A)(z^{k}-x^{\star})$
On the other hand, we can express the potential function as follows:
$\varphi_{\mu}(x,z)={}$ $\tfrac{1}{2}\langle(z-x^{\star}),A(z-x^{\star})\rangle+\tfrac{1}{2\mu}\|A(x-x^{\star})\|^{2}$
$={}$ $\tfrac{1}{2}\|A^{1/2}(z-x^{\star})\|^{2}+\tfrac{1}{2}\|\tfrac{1}{\sqrt{\mu}}A(x-x^{\star})\|^{2}$
$={}$ $\tfrac{1}{2}\Bigg\|\Bigg(\begin{array}[]{c}\tfrac{1}{\sqrt{\mu}}A(x-x^{\star})\\ A^{1/2}(z-x^{\star})\end{array}\Bigg)\Bigg\|^{2}.$ (76)
Using the AGD update, we can further write
$\Bigg(\begin{array}[]{c}\tfrac{1}{\sqrt{\mu}}A(x^{k+1}-x^{\star})\\ A^{1/2}(z^{k+1}-x^{\star})\end{array}\Bigg)={}$ $\Bigg(\begin{array}[]{c}\tfrac{1}{\sqrt{\mu}}A(I-\tfrac{1}{L}A)(\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}(x^{k}-x^{\star})+\tfrac{1}{\sqrt{\kappa}+1}(z^{k}-x^{\star}))\\ A^{1/2}[\tfrac{1}{\sqrt{\kappa}+1}(I-\tfrac{1}{\mu}A)(x^{k}-x^{\star})+\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}(I-\tfrac{1}{L}A)(z^{k}-x^{\star})]\end{array}\Bigg)=M\Bigg(\begin{array}[]{c}\tfrac{1}{\sqrt{\mu}}A(x^{k}-x^{\star})\\ A^{1/2}(z^{k}-x^{\star})\end{array}\Bigg)$ (83)
where $M=\Bigg(\begin{array}[]{cc}(I-\tfrac{1}{L}A)\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}&\tfrac{1}{\sqrt{\mu}}(A^{1/2}-\tfrac{1}{L}A^{3/2})\tfrac{1}{\sqrt{\kappa}+1}\\ \sqrt{\mu}(A^{-1/2}-\tfrac{1}{\mu}A^{1/2})\tfrac{1}{\sqrt{\kappa}+1}&(I-\tfrac{1}{L}A)\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}\end{array}\Bigg)$. Therefore,
$\varphi(x^{k+1},z^{k+1})=\tfrac{1}{2}\Bigg\|M\Bigg(\begin{array}[]{c}\tfrac{1}{\sqrt{\mu}}Ax^{k}\\ A^{1/2}z^{k}\end{array}\Bigg)\Bigg\|^{2}\leq\tfrac{1}{2}\|M^{\top}M\|\Bigg\|\Bigg(\begin{array}[]{c}\tfrac{1}{\sqrt{\mu}}Ax^{k}\\ A^{1/2}z^{k}\end{array}\Bigg)\Bigg\|^{2}=\|M^{\top}M\|\cdot\varphi(x^{k},z^{k})$
and it suffices to analyze $\|M^{\top}M\|$. Since $A\in\mathbb{S}^{n}_{++}$, without loss of generality, let $A=Q\Lambda^{2}Q^{\top}$ and
$M={}$ $\Bigg(\begin{array}[]{cc}(I-\tfrac{1}{L}A)\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}&\sqrt{\mu}(A^{-1/2}-\tfrac{1}{\mu}A^{1/2})\tfrac{1}{\sqrt{\kappa}+1}\\ \tfrac{1}{\sqrt{\mu}}(A^{1/2}-\tfrac{1}{L}A^{3/2})\tfrac{1}{\sqrt{\kappa}+1}&(I-\tfrac{1}{L}A)\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}\end{array}\Bigg)$ (86)
$={}$ $\Bigg(\begin{array}[]{cc}Q(I-\tfrac{1}{L}\Lambda^{2})Q^{\top}\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}&Q(\sqrt{\mu}\Lambda^{-1}-\tfrac{1}{\sqrt{\mu}}\Lambda)Q^{\top}\tfrac{1}{\sqrt{\kappa}+1}\\ \tfrac{1}{\sqrt{\mu}}Q(\Lambda-\tfrac{1}{L}\Lambda^{3})Q^{\top}\tfrac{1}{\sqrt{\kappa}+1}&Q(I-\tfrac{1}{L}\Lambda^{2})Q^{\top}\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}\end{array}\Bigg)$ (89)
$={}$ $\bigg(\begin{array}[]{cc}Q&\\ &Q\end{array}\bigg)\Bigg(\begin{array}[]{cc}(I-\tfrac{1}{L}\Lambda^{2})\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}&(\sqrt{\mu}\Lambda^{-1}-\tfrac{1}{\sqrt{\mu}}\Lambda)\tfrac{1}{\sqrt{\kappa}+1}\\ \tfrac{1}{\sqrt{\mu}}(\Lambda-\tfrac{1}{L}\Lambda^{3})\tfrac{1}{\sqrt{\kappa}+1}&(I-\tfrac{1}{L}\Lambda^{2})\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}\end{array}\Bigg)\bigg(\begin{array}[]{cc}Q&\\ &Q\end{array}\bigg)^{\top}$ (96)
$={}$ $\bigg(\begin{array}[]{cc}Q&\\ &Q\end{array}\bigg)N\bigg(\begin{array}[]{cc}Q&\\ &Q\end{array}\bigg)^{\top},$ (101)
where $N=\Bigg(\begin{array}[]{cc}(I-\tfrac{1}{L}\Lambda^{2})\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}&(\sqrt{\mu}\Lambda^{-1}-\tfrac{1}{\sqrt{\mu}}\Lambda)\tfrac{1}{\sqrt{\kappa}+1}\\ \tfrac{1}{\sqrt{\mu}}(\Lambda-\tfrac{1}{L}\Lambda^{3})\tfrac{1}{\sqrt{\kappa}+1}&(I-\tfrac{1}{L}\Lambda^{2})\tfrac{\sqrt{\kappa}}{\sqrt{\kappa}+1}\end{array}\Bigg)$ and it suffices to analyze $\|N^{\top}N\|=\|M^{\top}M\|$. By direct computation, $N^{\top}N$ is a 2 by 2 block matrix
$N^{\top}N=\frac{1}{(\sqrt{\kappa}+1)^{2}}\Bigg(\begin{array}[]{cc}\tfrac{L}{\mu}-\tfrac{1}{\mu}\Lambda^{2}-\tfrac{1}{L\mu}\Lambda^{4}+\tfrac{1}{\mu}\tfrac{1}{L^{2}}\Lambda^{6}&\sqrt{L}(\Lambda^{-1}-\tfrac{1}{L}\Lambda-\tfrac{1}{L\mu}\Lambda^{3}+\tfrac{1}{\mu L^{2}}\Lambda^{5})\\ \sqrt{L}(\Lambda^{-1}-\tfrac{1}{L}\Lambda-\tfrac{1}{L\mu}\Lambda^{3}+\tfrac{1}{\mu L^{2}}\Lambda^{5})&\mu\Lambda^{-2}+(\tfrac{L}{\mu}-2)I-\tfrac{1}{\mu}\Lambda^{2}+\tfrac{1}{L\mu}\Lambda^{4}\end{array}\Bigg),$
where each block is a diagonal matrix. Since permutation does not affect the spectrum, we can permute $N^{\top}N$ to obtain multiple 2 by 2 matrices, each taking form of
$\frac{1}{(\sqrt{\kappa}+1)^{2}}\Bigg(\begin{array}[]{cc}\tfrac{L}{\mu}-\tfrac{1}{\mu}\lambda^{2}-\tfrac{1}{L\mu}\lambda^{4}+\tfrac{1}{\mu L^{2}}\lambda^{6}&\sqrt{L}\lambda^{-1}-\tfrac{1}{\sqrt{L}}\lambda-\tfrac{1}{\sqrt{L}\mu}\lambda^{3}+\tfrac{1}{\mu L^{3/2}}\lambda^{5}\\ \sqrt{L}\lambda^{-1}-\tfrac{1}{\sqrt{L}}\lambda-\tfrac{1}{\sqrt{L}\mu}\lambda^{3}+\tfrac{1}{\mu L^{3/2}}\lambda^{5}&\mu\lambda^{-2}+\tfrac{L}{\mu}-2-\tfrac{1}{\mu}\lambda^{2}+\tfrac{1}{L\mu}\lambda^{4}\end{array}\Bigg),$
where $\lambda$ is one of the diagonal elements of $\Lambda$. Hence it suffices to compute
$\max_{\sqrt{\mu}\leq\lambda\leq\sqrt{L}}\Bigg\|\small{\Bigg(\begin{array}[]{cc}\tfrac{L}{\mu}-\tfrac{1}{\mu}\lambda^{2}-\tfrac{1}{L\mu}\lambda^{4}+\tfrac{1}{\mu L^{2}}\lambda^{6}&\sqrt{L}\lambda^{-1}-\tfrac{1}{\sqrt{L}}\lambda-\tfrac{1}{\sqrt{L}\mu}\lambda^{3}+\tfrac{1}{\mu L^{3/2}}\lambda^{5}\\ \sqrt{L}\lambda^{-1}-\tfrac{1}{\sqrt{L}}\lambda-\tfrac{1}{\sqrt{L}\mu}\lambda^{3}+\tfrac{1}{\mu L^{3/2}}\lambda^{5}&\mu\lambda^{-2}+\tfrac{L}{\mu}-2-\tfrac{1}{\mu}\lambda^{2}+\tfrac{1}{L\mu}\lambda^{4}\end{array}\Bigg)}\Bigg\|$ (104)
$={}$ $\max_{1\leq\lambda\leq\sqrt{\kappa}}\Bigg\|\small{\Bigg(\begin{array}[]{cc}\tfrac{L}{\mu}-\lambda^{2}-\tfrac{\mu}{L}\lambda^{4}+\tfrac{\mu^{2}}{L^{2}}\lambda^{6}&\sqrt{\frac{L}{\mu}}\lambda^{-1}-\sqrt{\frac{\mu}{L}}\lambda-\sqrt{\frac{\mu}{L}}\lambda^{3}+(\frac{\mu}{L})^{3/2}\lambda^{5}\\ \sqrt{\frac{L}{\mu}}\lambda^{-1}-\sqrt{\frac{\mu}{L}}\lambda-\sqrt{\frac{\mu}{L}}\lambda^{3}+(\frac{\mu}{L})^{3/2}\lambda^{5}&\lambda^{-2}+\tfrac{L}{\mu}-2-\lambda^{2}+\tfrac{\mu}{L}\lambda^{4}\end{array}\Bigg)}\Bigg\|$ (107)
$={}$ $\max_{1\leq\lambda\leq\sqrt{\kappa}}\Bigg\|\small{\Bigg(\begin{array}[]{cc}\kappa-\lambda^{2}-\tfrac{1}{\kappa}\lambda^{4}+\tfrac{1}{\kappa^{2}}\lambda^{6}&\sqrt{\kappa}\lambda^{-1}-\kappa^{-1/2}\lambda-\sqrt{\kappa^{-1/2}}\lambda^{3}+\kappa^{-3/2}\lambda^{5}\\ \sqrt{\kappa}\lambda^{-1}-\kappa^{-1/2}\lambda-\sqrt{\kappa^{-1/2}}\lambda^{3}+\kappa^{-3/2}\lambda^{5}&\lambda^{-2}+\kappa-2-\lambda^{2}+\kappa^{-1}\lambda^{4}\end{array}\Bigg)}\Bigg\|$ (110)
$={}$ $\max_{1\leq\lambda\leq\sqrt{\kappa}}\Bigg\|\small{\Bigg(\begin{array}[]{cc}\kappa-\lambda^{2}-\tfrac{1}{\kappa}\lambda^{4}+\tfrac{1}{\kappa^{2}}\lambda^{6}&\sqrt{\kappa}\lambda^{-1}-\kappa^{-1/2}\lambda-\sqrt{\kappa^{-1/2}}\lambda^{3}+\kappa^{-3/2}\lambda^{5}\\ \sqrt{\kappa}\lambda^{-1}-\kappa^{-1/2}\lambda-\sqrt{\kappa^{-1/2}}\lambda^{3}+\kappa^{-3/2}\lambda^{5}&\lambda^{-2}+\kappa-2-\lambda^{2}+\kappa^{-1}\lambda^{4}\end{array}\Bigg)}\Bigg\|$ (113)
$={}$ $\max_{1\leq\lambda\leq\sqrt{\kappa}}\Bigg\|\small{\Bigg(\begin{array}[]{cc}\frac{(\kappa-\lambda^{2})^{2}(\kappa+\lambda^{2})}{\kappa^{2}}&\frac{(\kappa-\lambda^{2})(\kappa-\lambda^{4})}{\kappa^{3/2}\lambda}\\ \frac{(\kappa-\lambda^{2})(\kappa-\lambda^{4})}{\kappa^{3/2}\lambda}&\frac{\lambda^{4}}{\kappa}+\kappa-\lambda^{2}+\frac{1}{\lambda^{2}}-2\end{array}\Bigg)}\Bigg\|$ (116)
Two eigenvalues of the matrix are respectively
$\tfrac{1}{2}(\tfrac{\lambda^{6}}{\kappa^{2}}\pm(\lambda^{4}-\kappa)\tfrac{\sqrt{4(\kappa-1)\kappa^{2}\lambda^{2}+\kappa^{2}+2\kappa(1-2\kappa)\lambda^{4}+\lambda^{8}}}{\kappa^{2}\lambda^{2}}+2\kappa-2\lambda^{2}+\tfrac{1}{\lambda^{2}}-2),$
whose maximizers are respectively $\sqrt{\kappa}$ and 1, giving
$\max\{-2+\tfrac{1}{\kappa}+\kappa,\tfrac{(\kappa-1)^{2}}{\kappa}(1+\tfrac{1}{2\kappa}+\tfrac{\sqrt{4\kappa+1}}{2\kappa})\}\leq\tfrac{(\kappa-1)^{2}}{\kappa}(1+\tfrac{1}{2\kappa}+\tfrac{\sqrt{4\kappa+1}}{2\kappa})$
and multiplying the upper bound by $\frac{1}{(\sqrt{\kappa}+1)^{2}}$ completes the proof.
Remark 4.
As a final remark, we compare the linear convergence rate for the new potential function with two standard convergence rates in the literature: $1-\frac{1}{\sqrt{\kappa}}$ for AGD and $(\frac{1-\sqrt{\kappa}}{1+\sqrt{\kappa}})^{2}$ for Chebeshev iteration [8]. As Figure˜6 (left) shows, our new convergence rate is sandwiched between these two rates. Figure˜6 (right) verifies the validity of this new potential function on a toy quadratic minimization problem.
Figure 6: Left: comparison between the convergence rate of the new potential function and classic ones . Right: plots of the new potential function, the classic potential function, and the function value gap
C.3 Details of in Subsection 5.4
In this section, we illustrate the convergence analysis of a variant of OSGM-H. We make the following assumptions throughout the section.
  1. A1:
    $f$ is $L$-smooth and convex with $\operatorname{dom}f=\mathbb{R}^{n}$.
  2. A2:
    The proximal term $w$ is convex.
Similarly, we denote $\varphi^{\star}\coloneqq\varphi(x^{\star})$. The challenging part of the analysis is that the proximal term $w$ will introduce non-Lipschitzness of the feedback function if $\|\mathcal{G}_{L}(x)\|\rightarrow 0$. To this end, we consider Algorithm˜8.
1input initial point $x^{1}$, initial stepsize $P_{1}\in\mathbb{R}^{n\times n}$, online proximal regularizer $\eta>0$
2for $k=1,2,\dots$ do
3 $P_{k+1}={}\operatornamewithlimits{arg\,min}_{P}\big\{\tfrac{f(x^{k}-P_{k}\mathcal{G}_{L}(x^{k}))+\langle\nabla f(x^{k}-P_{k}\mathcal{G}_{L}(x^{k}))\mathcal{G}_{L}(x^{k})^{\top},P-P_{k}\rangle+\omega(x^{k}-P\mathcal{G}_{L}(x^{k}))}{\|\mathcal{G}_{L}(x)\|^{2}}+\tfrac{1}{2\eta}\|P-P_{k}\|_{F}^{2}\big\}$
4  Choose $x^{k+1}$ such that $\varphi(x^{k+1})\leq{}\min\{\varphi(x^{k}-P_{k+1}\mathcal{G}_{L}(x^{k})),\varphi(x^{k})\}$
5 end for
Algorithm 8 OSGM-H with proximal gradient update
This update corresponds to Classic-HDM we discussed in Section˜2.2. As a first step, we define progress $h_{k}\coloneqq\tfrac{\varphi(x^{k+1})-\varphi(x^{k})}{\|\mathcal{G}_{L}(x^{k})\|^{2}}$. Note that $h_{k}\leq 0$ by definition, which enables the following progress reduction.
Lemma C.3 (Nonsmooth reduction).
Under A1 and A2, $\min_{1\leq k\leq K}\|\mathcal{G}_{L}(x^{k})\|^{2}\leq\tfrac{\varphi(x^{1})-\varphi^{\star}}{K}\tfrac{1}{\frac{1}{K}\sum_{k=1}^{K}-h_{k}}$.
Proof.
The proof resembles that of the hypergradient reduction.
$(\textstyle\sum_{k=1}^{K}-h_{k})\cdot\displaystyle\min_{1\leq k\leq K}\|\mathcal{G}_{L}(x^{k})\|^{2}\leq{}$ $\textstyle\sum_{k=1}^{K}-h_{k}\|\mathcal{G}_{L}(x^{k})\|^{2}$ (by $h_{k}\geq 0$)
$={}$ $-[\textstyle\sum_{k=1}^{K}\varphi(x^{k+1})-\varphi(x^{k})]$ (by definition of $h_{k}$)
$={}$ $\varphi(x^{1})-\varphi(x^{K+1})$
$\leq{}$ $\varphi(x^{1})-\varphi^{\star}.$ (by $f(x^{K+1})\leq f^{\star}$)
An important step is to establish online learning guarantees with respect to $\sum_{k=1}^{K}h_{x^{k}}(P_{k+1})$. Below, we prove the result for a more general setting.
Lemma C.4 (Prescient online proximal gradient).
Suppose $\{(\ell_{k},\pi_{k})\}$ are a sequence of convex functions and $\{\ell_{k}\}$ are $L$-smooth. Consider the online proximal gradient update
$P_{k+1}=\operatornamewithlimits{arg\,min}_{P\in\mathcal{P}}~\{\ell_{k}(P_{k})+\langle\nabla\ell_{k}(P_{k}),P-P_{k}\rangle+\pi_{k}(P)+\tfrac{1}{2\eta}\|P-P_{k}\|^{2}_{F}\}.$
Then with $\eta\leq\frac{1}{L}$, $\sum_{k=1}^{K}\ell_{k}(P_{k+1})+\pi_{k}(P_{k+1})-[\ell_{k}(\hat{P})+\pi_{k}(\hat{P})]\leq\frac{1}{2\eta}\|P_{1}-\hat{P}\|^{2}_{F}$ for any $\hat{P}\in\mathcal{P}$.
Proof.
By the optimality condition of the online update, for any $\hat{P}\in\mathcal{P}$,
$\langle\nabla\ell_{k}(P_{k}),P_{k+1}-P_{k}\rangle+\pi_{k}(P_{k+1})+\tfrac{1}{2\eta}\|P_{k+1}-P_{k}\|^{2}_{F}$
$\leq{}$ $\langle\nabla\ell_{k}(P_{k}),\hat{P}-P_{k}\rangle+\pi_{k}(\hat{P})+\tfrac{1}{2\eta}\|\hat{P}-P_{k}\|^{2}_{F}-\tfrac{1}{2\eta}\|P_{k+1}-\hat{P}\|^{2}_{F}.$
Re-arranging the terms,
$\|P_{k+1}-\hat{P}\|^{2}_{F}\leq{}$ $\|P_{k}-\hat{P}\|^{2}_{F}-\|P_{k+1}-P_{k}\|^{2}_{F}+2\eta\langle\nabla\ell_{k}(P_{k}),\hat{P}-P_{k+1}\rangle+2\eta[\pi_{k}(\hat{P})-\pi_{k}(P_{k+1})]$
$={}$ $\|P_{k}-\hat{P}\|^{2}_{F}-\|P_{k+1}-P_{k}\|^{2}_{F}+2\eta\langle\nabla\ell_{k}(P_{k}),\hat{P}-P_{k}\rangle$
$\quad+2\eta\langle\nabla\ell_{k}(P_{k}),P_{k}-P_{k+1}\rangle+2\eta[\pi_{k}(\hat{P})-\pi_{k}(P_{k+1})]$
[…]
$\quad+2\eta\langle\nabla\ell_{k}(P_{k}),P_{k}-P_{k+1}\rangle+2\eta[\pi_{k}(\hat{P})-\pi_{k}(P_{k+1})]$ (Convexity of $\ell_{k}$)
[…]
$+2\eta[\ell_{k}(P_{k+1})-\ell_{k}(P_{k})+\langle\nabla\ell_{k}(P_{k}),P_{k}-P_{k+1}\rangle-\tfrac{1}{2\eta}\|P_{k+1}-P_{k}\|^{2}_{F}].$
When $\eta\leq\tfrac{1}{L}$, we have, by smoothness of $\ell_{k}$, that
$\ell_{k}(P_{k+1})-[\ell_{k}(P_{k})+\langle\nabla\ell_{k}(P_{k}),P_{k+1}-P_{k}\rangle]\leq\tfrac{L}{2}\|P_{k+1}-P_{k}\|^{2}_{F}\leq\tfrac{1}{2\eta}\|P_{k+1}-P_{k}\|^{2}_{F}$
[…]
Chaining Lemma˜C.3, Lemma˜C.4 and the fact $h_{x}(\frac{1}{L}I)\leq-\frac{1}{2L}$ [3, Corollary 10.18] gives the desired result.
Theorem C.1 (Theorem˜5.2).
Suppose $f$ is $L$-smooth convex and $w$ is convex. Then for any benchmark stepsize $\hat{P}\in\mathcal{P}$, Algorithm˜8 achieves
$\min_{1\leq k\leq K}\|\mathcal{G}_{L}(x^{k})\|^{2}\leq\tfrac{\varphi(x^{1})-\varphi^{\star}}{K}\tfrac{1}{\frac{1}{K}\max\{\sum_{k=1}^{K}h_{x^{k}}(\hat{P})-\frac{L}{2}\|P_{1}-\hat{P}\|_{F}^{2},0\}}.$
In particular, taking $P_{1}=\hat{P}=\frac{1}{L}I$ yields $\min_{1\leq k\leq K}\|\mathcal{G}_{L}(x^{k})\|^{2}\leq\tfrac{2L[\varphi(x^{1})-\varphi(x^{\star})]}{K}$.
Proof.
Using Lemma˜C.4, we can bound
$\sum_{k=1}^{K}h_{k}\leq\sum_{k=1}^{K}h_{x^{k}}({P_{k+1}})\leq\sum_{k=1}^{K}h_{x^{k}}(\hat{P})+\tfrac{1}{2\eta}\|P_{1}-\hat{P}\|_{F}^{2}.$
Plugging the relation back into Lemma˜C.3 and using $h_{x^{k}}(\tfrac{1}{L}I)\leq-\tfrac{1}{2L}$ completes the proof. ∎